### v1

Abstract
Dynamic resource management is an increasingly important capability of High Performance Computing systems, as it enables jobs to adjust their resource allocation at runtime. This capability has been shown to reduce workload makespan, substantially decrease job waiting times and improve overall system utilization. In this context, malleability refers to the ability of applications to adapt to new resource allocations during execution. Although beneficial, malleability incurs significant reconfiguration costs, making the reduction of these costs an important research topic.
Some existing methods for MPI applications respawn the entire application, which is an expensive solution that avoids the reuse of original processes. Other MPI methods reuse them, but fail to fully release unneeded processes when shrinking, since some ranks within the same communicator remain active across nodes, preventing the application from returning those nodes to the system. This work overcomes both limitations by proposing a novel parallel spawning strategy, in which all processes cooperate in spawning before redistribution, thereby reducing execution time. Additionally, it removes shrinkage limitations, allowing better adaptation of parallel systems to workload and reducing their makespan. As a result, it preserves competitive expansion times with at most a $1.25\times$ overhead, while enabling fast shrink operations that reduce their cost by at least $20\times$. This strategy has been validated on both homogeneous and heterogeneous systems and can also be applied in shared-resource environments.
keywords:
[…]
1 Introduction
Despite advances in High Performance Computing (HPC) systems during the exascale era, the efficient utilisation of resources such as CPUs, GPUs and memory remains a significant challenge [12]. Scientific applications often underuse their allocations due to factors such as overcommitment, bottlenecks or limited scalability.
To address this issue, Resource Management Systems (RMS) must be able to distribute resources dynamically, a capability known as Dynamic Resource Management (DRM). This capability enables RMSs to adjust job allocations at runtime, provided the applications are malleable, that is, capable of resizing and adapting to changes in resource availability during execution. Such flexibility allows applications to achieve more effective optimization of resource utilization [4], throughput [9], energy efficiency [3, 11], and I/O performance [23].
Recent studies by the US Exascale Computing Project [2] suggest that DRM could significantly enhance the performance of both applications and systems, making it a key contribution for improving relevant metrics in future platforms [13]. However, its adoption remains limited compared to other MPI features [6], primarily due to the complexity involved in adapting applications to support malleability and enabling systems to dynamically reallocate resources at runtime.
There are already libraries available to perform reconfigurations in applications; some of the most relevant are DMR[9], FlexMPI[14], MaM[16] or DPP[8]. A particularly challenging scenario for most of these tools is shrinking, i.e., reducing the number of active processes in a job at runtime to free resources or adapt to changing workload demands. This operation is limited by the rigid structure of MPI communicators, especially the standard MPI_COMM_WORLD (MCW), which does not allow the partial termination of ranks.
As shown in [20], terminate-based (TB) methods can shrink without spawning new processes, providing a time efficient alternative to traditional spawn-based (SB) approaches. Nevertheless, a major limitation of TB shrinking is the persistence of zombie processes, ranks that remain asleep until the application terminates, as they cannot be fully removed from the global MCW. This creates an issue: the nodes on which these zombies are mapped cannot be returned to the RMS until they terminate, preventing node reuse. A common workaround involves isolating each MCW communicator on a separate node, allowing finer control over process groups. However, this solution depends on node-specific spawning, which suffers from poor scalability due to its inherently sequential nature [22].
To overcome these constraints, we propose a novel technique that retains the benefits of TB shrinking while addressing its key limitations, supporting both expansion and shrinkage at minimal cost. Our approach isolates each MCW communicator on a separate node and enables all available processes to perform spawning in parallel rather than sequentially, thereby reducing the cost of expansion. In addition, it avoids spawning new processes during shrinkage and eliminates the persistence of zombie processes while maintaining the high performance observed in TB methods. These improvements make our solution particularly suitable for DRM in large-scale HPC environments, with full support for heterogeneous systems and shared resources. The approach has been included in the MaM library [16], which integrates multiple techniques for resizing and data redistribution, allowing the selection of the optimal solution depending on the context.
The main contributions of this work are:
  • The design and implementation of two novel parallel spawning algorithms tailored to the MPI model, enabling scalable malleability for HPC applications.
  • The extension of the MaM library [16] to support dynamic reconfiguration in heterogeneous queue systems and shared-resource environments.
  • A comprehensive performance evaluation of the proposed algorithms on both homogeneous and heterogeneous clusters, demonstrating their advantages over the previous best method in terms of scalability and reconfiguration overhead.
The remainder of this paper is organized as follows: Section 2 presents a comprehensive review of related work on MPI malleability and spawning techniques. Section 3 provides an overview of process spawning mechanisms in MPI and their integration into MaM. Section 4 explains the two new parallel spawning algorithms proposed in this work. Section 5 presents an evaluation of their performance. Finally, Section 6 concludes the paper and outlines directions for future work.
2 Background
According to the classification proposed by Feitelson and Rudolph [5], applications executed in large-scale computing facilities can be grouped into four categories. This classification is based on two main criteria: (1) whether the number of processes can change during execution, i.e., whether reconfiguration is supported, and (2) who determines the job size, either the user or the RMS. Table 1 summarizes this classification.
Job Type Resource Allocation Who Sets Size Description
Rigid Job Static User Can only be executed with a fixed number of processes defined by the user. No reconfiguration is allowed.
Moldable Job Static RMS Can start with a variable number of processes. The RMS determines the size when the execution begins.
Evolving Job Dynamic User Includes a user-defined reconfiguration scheme. The RMS must meet the resource requests for the job to continue executing.
Malleable Job Dynamic RMS Can be reconfigured at runtime if the RMS decides to adjust the resource allocation.
Table 1: Classification of parallel jobs according to Feitelson and Rudolph [5].
In this paper, we consider malleability as the ability of a distributed parallel job to change its size in terms of MPI ranks by adjusting the computational resources initially allocated to the job at any point during execution as often as required. This capability is activated at specific checkpoints within the application and requires the execution of a series of stages:
  1. 1.
    Reconfiguration feasibility: The RMS determines whether the job should be resized according to a dynamic resource allocation policy. If not, the subsequent steps are not performed.
  2. 2.
    Process management. The RMS assigns a new number of cores to the job, which dictates how many MPI ranks should be created or terminated. Processes before resizing are considered sources, while those that continue after resizing are considered targets.
  3. 3.
    Data redistribution: Sources transfer their data to targets.
  4. 4.
    Resume execution. The application resumes execution with the targets.
Incorporating all these stages into a parallel application is a complex task that goes far beyond simple integration with the RMS. It requires modifications to the application’s structure to support dynamic creation and termination of processes, manage data redistribution efficiently, and reconfigure communication patterns on the fly. These challenges make dynamic applications difficult to implement from scratch, often requiring dedicated tools and frameworks that simplify the transformation of traditional applications into dynamic ones. Without such support, adapting applications to handle resizing becomes both error-prone and time-consuming.
One malleable solution is Flex-MPI [14], a library built on top of MPICH [21] that bases its spawn operations on the MPI_Comm_spawn function. Its primary goal is to help applications become dynamic and the main result of this solution is the Merge method, which is described below. When an application expands from $x$ processes to $y$ processes, only the additional yxy-x processes are spawned and added to the existing ones. Then, each new process is created through a separate call to the spawn function, which allows a fine grain when shrinking the application, as dynamic processes can be terminated at any time. Nevertheless, there is a limitation with the minimum number of ranks: the initial number of ranks cannot be terminated until the end of the application.
A similar solution is MaM [16], a malleability module also built on top of MPICH and within the MPI scope, developed to analyze and compare different malleability techniques. Like Flex-MPI, it bases its spawn operation on the MPI_Comm_spawn function, providing a Merge method, but without the limitation of the minimum number of processes. MaM is part of the Proteo framework [19], which helps transform applications into malleable ones, supports the generation of dynamic workloads, and enables performance analysis of different resize operations to select the optimal alternative for a given situation.
Another approach is DPP [8], which explores the new concept of MPI Sessions, extending it to implement on-the-fly dynamic mechanisms. Since the sessions model operates outside the traditional MPI environment, the MCW communicator does not exist and some functions, such as MPI_Comm_spawn, cannot be used. However, alternative non-standard MPI functions are provided to allow communication between ranks. To spawn processes, it uses a custom dynamic RMS, which determines whether a job should get or return resources. When a job gets additional resources, it informs a PMIx Reference Runtime Environment (PRRTE) daemon, which then spawns of new processes. Once the processes are ready, the job is informed of the new resources and how to connect with them, which is managed by DPP. During shrink, processes do not have the limitations of the traditional environment and can be terminated normally.
Lastly, an alternative solution is Dynamic Management of Resources library (DMRlib) [9], which consists of two components: a parallel distributed runtime based on MPI and an extension of SLURM to enable the execution of malleable jobs. Initially, the library only allowed the simplest approach, the Baseline method, which always spawns the new number of processes $y$ and terminates the previous $x$ processes. However, the library now supports other environments for process management, such as DPP [7] or MaM [10], providing a wider range of possibilities.
For further details on MPI spawn based malleability, see [1], while for a broader review of dynamic resource management [24] is recommended.
3 MaM overview
MaM is the Proteo module responsible for completing malleability. It provides a set of features designed to transform traditional MPI applications into malleable ones, requiring minimal effort from developers. Additionally, it defines a simple and intuitive interface, which lowers the barrier to integrating malleability into existing parallel applications [16].
Specifically, MaM implements multiple methods and strategies to handle the most critical stages of malleability, process management and data redistribution, allowing applications to dynamically adjust their number of MPI ranks during execution.
For the process management stage, i.e., resizing from an initial group of NS processes to a new group of NT processes, MaM offers two main approaches: the Baseline method, which always creates a new group of processes, and the Merge method, which reuses existing processes and only spawns or terminates the difference in ranks. These methods can be combined with two strategies: the Asynchronous strategy, which overlaps process creation with application execution to reduce downtime, and the Single strategy, where only one process performs the spawn operation and informs the rest afterwards. This flexible design allows users to decide the preferred configuration according to their application needs.
For the data redistribution stage, MaM distinguishes between constant and variable data, as this affects how and when the redistribution can take place. While variable data requires execution to be paused until the transfer is complete, constant data can be redistributed asynchronously, allowing overlap with ongoing computation. The library also supports three main communication methods for this stage: Point-to-Point, Collective, or One Sided primitives.
Additionally, asynchronous redistribution can be implemented through two strategies: Non-blocking communication, which relies on routines like MPI_Test to monitor progress, and Threading, where dedicated threads handle data transfer independently of the main computation.
Using MaM to incorporate malleability into an application requires specifying a method with none, one, or more of the strategies for stages 22 and $3$. A more detailed description of all alternatives related to process management can be found in [20] and to automatic data redistribution in [15] and [18].
4 Methods for spawning processes in parallel
As previously discussed, the MaM shrinkage approach is constrained by its inability to fully terminate unneeded processes during execution. As a consequence, is not possible to return those resources until the MCW they pertain terminates completely. To overcome these issues, this work proposes new strategies based on parallel spawning, which allows processes to be resized efficiently and overcome said situation.
Parallel process creation is inherently complex and is coordinated across multiple phases: (1) It begins with the spawn of new process groups and opening communication ports. (2) it is followed by a synchronization phase, where all groups wait until every required port is ready. (3) Next, the newly spawned groups are merged into a single communicator through the ports. (4) Finally, a global reordering of ranks is performed to maintain a logical order between nodes.
The first phase has been examined through two alternatives: A simpler option only for homogenous systems called Hypercube strategy, and an improved but more complex one that works for heterogenous or shared systems, called Iterative Diffusive strategy. The following sections detail each of these steps, highlighting the specific operations involved.
4.1 Hypercube: Parallel Spawning for Homogeneous Systems
This parallel strategy is based on performing a call to the MPI_Comm_spawn function for each node that is to be used, assigning all available cores of the node to each newly created process.
The procedure is organized into a series of steps. In each step, all active processes, whether source or target, concurrently execute a spawn operation, each one acting on a different node. As the number of target processes increases, so does the number of processes participating in the spawning operation. Consequently, with each successive step, a growing number of processes are involved in initiating new spawns, enabling a scalable and distributed expansion across nodes.
In the first step, only the source processes participate in the operation. In total, up to $C\cdot I$ processes can be launched, where $C$ is the number of cores per node and $I$ is the number of initial nodes. From the second step onward, both the original and previously created processes participate, exponentially increasing the spawning capacity. However, in the final step, only the first $x$ processes participate, where $x$ corresponds to the remaining number of nodes needed to reach the desired number of nodes.
As an example, consider a system with 20 cores per node. Starting with a single fully occupied node, the first step could launch up to 20 additional nodes. In the second step, with 420 processes available (21 nodes with 20 cores each), it would be possible to occupy 410 additional nodes by creating 20 processes per node.
From a technical standpoint, this strategy is implemented using MPI_Comm_spawn in combination with the MPI_COMM_SELF communicator and an MPI_Info object that specifies the target node for each new group of processes. Furthermore, all processes in each node are assigned the group identifier ranging from 0 to the total number of groups to be created (group_id), identifying which group it belongs.
The number of steps $s$ required to reach a total of $N$ nodes, starting from $I$ initial nodes and with $C$ cores per node, can be calculated using Equation 1. $N$ and $I$ are calculated as $\frac{NT}{C}$ and $\frac{NS}{C}$, respectively, where $NT$ is the amount of targets, and $NS$ the amount of sources.
${s}=\frac{\ln(N/I)}{\ln(1+C)}$ (1)
From this, the total number of nodes $T_{s}$ at the end of step $s$ can be computed as:
${T_{s}}=(1+C)^{s}\cdot I$ (2)
And the corresponding number of processes spawned $t_{s}$ in step $s$ is simply:
${t_{s}}=T_{s}\cdot C$ (3)
Figure 1 illustrates this strategy with a simple example in which the number of processes is expanded from $NS=1$ to $NT=8$, assuming an architecture with $C=1$ core per node. In this case, seven additional groups are created over three steps. The process is visualized as a cube, where the edges indicate which group spawns which in each step, and the vertices are labelled with the corresponding group_id, or with the label $I$ for the initial group. By the end of the procedure, a total of eight nodes are occupied.
Figure 1: Generation of 7 groups using parallel spawning. Each step is represented with different markers.
4.2 Iterative Diffusive: Recursive Algorithm for Heterogeneous Systems
This method improves upon the previous scheme by including the ability to create process groups of variable sizes at each step, which the Hypercube is not capable of. This makes it particularly well-suited to heterogeneous or shared computing environments, where nodes may have different capacities or resources may be distributed unevenly.
As in the original scheme, the algorithm proceeds in discrete steps, where all existing processes collaborate to spawn new groups on additional nodes. The assignment of a unique group identifier (group_idgroup\_id), ranging from 0 to the total number of groups to be created, is also preserved.
Three vectors are needed to describe the system configuration. The number of entries in these vectors is equal to the number of nodes in the system (N).
  • $M$: Contains the number of available cores assigned to a job on each node.
  • AA: Stores the number of processes running on each node before starting this algorithm.
  • $S$: Has the number of processes to be created in each node when executing this algorithm.
Based on these vectors, a stepwise sequence can be constructed to determine the total number of processes and nodes involved at each spawning step. Only nodes with positive entries in the vector $S$ will allocate new processes; thus, each non-zero entry indicates the creation of a new group on the corresponding node. This vector is computed from the previous ones as follows:
$S_{i}=M_{i}-A_{i},\quad\text{for }i=0,1,\dots,N-1$
Equation 4 defines the total number of processes $t_{s}$ present in step $s$, as a function of the cumulative total of previous steps ($t_{s-1}$) and the number of processes generated in the current step ($g_{s}$). Step $s=0$ corresponds to the initial state, which includes only the source processes, as specified by the vector $A$.
$t_{s}=\begin{cases}\sum\limits_{j=0}^{N-1}A_{j},&\text{if }s=0\\ t_{s-1}+g_{s},&\text{if }s>0\end{cases}$ (4)
Equation 5 defines the number of processes $g_{s}$ generated in step $s$. At each stage, the elements of vector $S$ from index $\nu_{s}$ to index $\lambda_{s}$ are accumulated to calculate the number of new processes that need to be generated.
$g_{s}$ $=\sum\limits_{j=\nu_{s}}^{\lambda_{s}}S_{j}$ $\quad\text{where}\quad\lambda_{s}$ $=\min\left(N-1,\,\left(\sum\limits_{i=0}^{s-1}t_{i}\right)-1\right),$ (5)
$\nu_{s}$ $=\sum\limits_{i=0}^{s-2}t_{i}$
Finally, Equation 7 represents the cumulative number of occupied nodes ($T_{s}$) up to step $s$, whereas Equation 6 specifies the number of new nodes ($G_{s}$) occupied by newly created groups during that step. In all equations, null elements in vector $S$ are disregarded.
$G_{s}=\min(N-T_{s-1},\,t_{s-1})$ (6)
$T_{s}=\begin{cases}I,&\text{if }s=0\\ T_{s-1}+G_{s},&\text{if }s>0\end{cases}$ (7)
4.3 Synchronization Between Process Groups
Once all processes have been created, it is necessary to establish connections between them to enable direct communication. With the adopted creation strategy, each newly spawned group initially knows only its parent process (the one that created it) and, if applicable, its own child processes. As shown in Figure 1, processes can only communicate directly with those to which they are connected by an edge in the spawn graph.
To enable communication between separate groups, the functions MPI_Comm_connect and MPI_Comm_accept are used. These functions establish an inter-communicator between two independent groups. However, to use them correctly, certain preliminary steps must be taken.
First, the process groups must be prepared for connection. The processes designated as roots in the MPI_Comm_accept calls must first open a port using MPI_Open_port and publish the service name via MPI_Publish_name. Since MPI_Comm_connect must be called after the corresponding port has been opened by the other group, a global synchronization phase between all groups is required. Otherwise, execution errors may occur, as observed with implementations such as MPICH [^1] This behaviour has only been tested with MPICH..
The synchronization between groups consists of three steps and is carried out within each group using a dedicated subcommunicator. The objective is to ensure that all processes are aware of the port status before attempting to establish a connection.
  1. [1]
  2. 1.
    Subcommunicator creation. Each group creates a subcommunicator using MPI_Comm_split, including the root process of the group and all processes that have created child groups. This subcommunicator will be used to coordinate intra-group synchronization.
  3. 2.
    Upside Synchronization. Each process that has spawned child groups waits to receive a token from each of them using MPI_Irecv followed by MPI_Waitall. Then, all processes in the subcommunicator execute a call to MPI_Barrier, ensuring that all tokens have been received. Finally, the root process (if not part of the source group) notifies its parent group using MPI_Send.
  4. 3.
    Downside Synchronization The root process of each group (except the source group) waits to receive a token from its creator via MPI_Recv. Then, processes in the subcommunicator perform a MPI_Barrier call (excluding the source group). Finally, each process that has spawned child groups sends them a token using MPI_Isend followed by MPI_Waitall. These barriers are essential to ensure that no group proceeds before all of its descendants have completed the Upside step, and to prevent any group from notifying its children before being notified by its parent group.
Listing 1 shows the implementation of the synchronization mechanism. The processes use the world_c communicator, which corresponds to MCW for spawned processes, or to a shared communicator among source processes in the case of creators. The parent_c communicator is an intercommunicator between a spawned group and its parent. The variable qty_s indicates the number of child groups spawned by a process, and the corresponding intercommunicators are stored in the spawn_c vector.
1void common_synch(MPI_Comm world_c, MPI_Comm parent_c,
2 int qty_s, MPI_Comm *spawn_c) {
3 int myId, i, root = 0;
4 int color = qty_s ? 1 : MPI_UNDEFINED;
5 MPI_Request *reqs = malloc(qty_s * sizeof *reqs);
6 MPI_Comm synch_ranks;
7
8 //Only root ranks or ranks with children synchronize
9 MPI_Comm_rank(world_c, &myId);
10 MPI_Comm_split(world_c, color, myId, &synch_ranks);
11
12 // Wait for all children of this group
13 for(i=0; i<qty_s; i++) {
14 MPI_Irecv(…, source=root, spawn_c[i], &reqs[i]);
15 }
16 if(qty_s){MPI_Waitall(reqs);MPI_Barrier(synch_ranks);}
17
18 if(parent_c != MPI_COMM_NULL && myId == root) {
19 // Send to parent this group is ready
20 MPI_Send(…, dest=root, parent_c);
21 // Wait for the other groups readiness
22 MPI_Recv(…, source=root, parent_c);
23 }
24
25 // Notify this group that all groups are ready
26 if(parent_c != MPI_COMM_NULL && qty_s)
27 {MPI_Barrier(synch_ranks);}
28 // Notify children that all groups are ready
29 for(i=0; i<qty_s; i++)
30 {MPI_Isend(…, dest=root, spawn_c[i], &reqs[i]);}
31 if(qty_s) {MPI_Waitall(reqs);}
32
33 if(synch_ranks != MPI_COMM_NULL)
34 {MPI_Comm_disconnect(&synch_ranks);}
35}
Listing 1: Syncrhonization step before binary connection.
Figure 2 illustrates the upside synchronization and downside synchronization steps, shown on the left and right sides, respectively. In this example, six process groups are created in two spawning steps. The arrows indicate which processes send messages to signal their readiness in the upside phase, and to confirm that all groups are ready in the downside phase. A group represented in black with stripes indicates that it must wait to receive all messages before proceeding. At that point, a MPI_Barrier is used to synchronize the group, after which the execution can continue.
Figure 2: Synchronization of 6 spawned groups and the initial node. The arrows indicate the direction of messages, and stripes the need of that group of using a Barrier.
4.4 Binary connection
Once synchronization between all groups is complete, the connection operation can begin. The adopted strategy consists of making pairwise binary connections in several successive steps. In each step, groups with an identifier less than half the number of active groups perform an MPI_Comm_accept operation, while those with a higher identifier execute an MPI_Comm_connect. In the case of an odd number of groups, the middle group is temporarily excluded from that step and proceeds directly to the next one.
Each group connects to the group whose identifier is its mirror image within the current set of active groups, that is, the group symmetrically positioned relative to the center. Once the connection is completed, both groups are merged into a new communicator and adopt a common group identifier, corresponding to the group that executed the MPI_Comm_accept call.
This procedure is repeated, halving the number of active groups at each step, until all processes are connected within a single group. Figure 3 shows an example in which seven spawned groups are connected in three steps, continuing the scenario previously illustrated in Figure 1.
Thanks to the synchronization mechanism described earlier, no strict ordering is required for the execution of Connect operations. Groups performing accept simply wait for incoming connections, regardless of the order in which they arrive.
Figure 3: Connection of 7 spawned groups in 3 steps. In each step the amount of groups is halved and connections operations may be unordered.
Listing 2 presents a simplified version of the code used to implement this binary connection strategy between groups. In the code, the groups are distinguished based on whether they should perform a MPI_Comm_accept operation (lines L11–L15) or a MPI_Comm_connect (lines L17–L23). In the latter case, the group obtains the corresponding port name (line L19) using the group_id of the target group and the MPI_Lookup_name function.
1void binary_connection(int groups, int group_id,
2 char *my_port, MPI_Comm *new_comm) {
3 int middle, new_groups, new_gid;
4 char *port;
5 MPI_Comm merge_comm, aux_comm, intercomm;
6 merge_comm = aux_comm = MPI_COMM_WORLD;
7 intercomm = MPI_COMM_NULL;
8 while(groups > 1) {
9 middle = groups / 2;
10 new_groups = ceil(groups / 2.0);
11 if(group_id < middle) {
12 MPI_Comm_accept(my_port, MPI_INFO_NULL,
13 MAM_ROOT, merge_comm, &intercomm);
14 MPI_Intercomm_merge(intercomm, 0, &aux_comm);
15 merge_comm = aux_comm;
16
17 } else if(group_id >= new_groups) {
18 new_gid = groups - group_id - 1;
19 get_remote_port(new_gid, &port);
20 MPI_Comm_connect(port, MPI_INFO_NULL,
21 MAM_ROOT, merge_comm, &intercomm);
22 MPI_Intercomm_merge(intercomm, 1, &aux_comm);
23 merge_comm = aux_comm; group_id = new_gid;
24
25 } groups = new_groups;
26 } *new_comm = merge_comm;
27}
Listing 2: Binnary connection step for the parallel spawn.
4.5 Rank Reordering and Connection with sources
Once all the generated groups have been merged into a single one, the process ranks may not be ordered across nodes. This is because binary connections between groups do not enforce a specific order and are susceptible to race conditions.
To ensure a consistent global order, an MPI_Comm_split call is made, where each process specifies its new rank as follows:
$\text{world\_rank}+\sum\limits_{j=0}^{N-1}A_{j}+\sum\limits_{j=0}^{\text{group\_id-1}}S_{j}$
At the same time, all intermediate communicators created throughout the strategy are disconnected using MPI_Comm_disconnect. This frees up resources and ensures that any node can be removed or replaced in future reconfigurations without leaving residual dependencies.
Finally, the target group is connected to the source group, establishing the final inter-communicator that links them. This operation concludes the process creation and connection phases.
5 Experimental results
5.1 Hardware and Software
The experiments were conducted on two clusters. The first one, MarenostrumV [^2] https://www.bsc.es/supportkc/docs/MareNostrum5/overview (MNV), used 32 nodes from the general queue. Each node has two 56-core Intel Xeon 8480 CPUs (3584 cores in total). This system was employed to analyze the behaviour of both spawning techniques described in Section 4.
The second cluster, Nasp [^3] https://www.hpca.uji.es/computing-facilities/, consists of two sets of 8 nodes. The first set has nodes with two 10-core Intel Xeon 4210 CPUs (160 cores total), connected through a 100 Gbps InfiniBand EDR and a 10 Gbps Ethernet network. The second set uses nodes with 32-core Intel Xeon 6346 CPUs (256 cores total) connected via 10 Gbps Ethernet. Both sets communicate through a shared 10 Gbps Ethernet link between their switches. This heterogenous cluster was used to evaluate the effectiveness of the improved method in Section 4.2.
All experiments in MNV were run with MPICH 4.2.0 and CH4:OFI (InfiniBand), while in NASP with MPICH 3.4.3 and CH3:Nemesis (Ethernet). Both the Proteo version [^4] https://lorca.act.uji.es/gitlab/martini/malleability_benchmark/-/tree/Paper-ParallelSpawn and the experimental results [17] are publicly available.
Proteo was used to execute the different approaches. Before reconfiguration, 5 iterations of Pi computation with MPI_Allgather were performed to ensure MPI initialization. MaM was configured with Baseline (B) and Merge (M), each with both strategies described in 4, or Merge without strategies. For shrinking, only Baseline plus the described techniques was compared against a stand-alone Merge. Each configuration was executed 20 times, and the median result was reported for every triplet of sources, targets, and method.
5.2 Homogenous System Evaluation
On the homogeneous cluster MNV, both strategies described in Section 4 were evaluated. In total, 5 configurations were tested for expansion and 3 for shrinking. Each configuration involved a single reconfiguration from $NS$ to $NT$ processes, using up to 42 combinations of nodes from the set 1,2,4,8,16,24,32{1,2,4,8,16,24,32}.
Figure 4 reports the median reconfiguration times depending on the number of sources ($NS$) and targets ($NT$). The top Figure 4(a) shows expansion results, where the Merge method outperforms all others in 17 out of 21 cases ($80.9\%$). Parallel Merge follows closely, requiring up to $1.13\times$ more time due to additional synchronizations among groups. This overhead grows when more than 8 groups are created and the number of groups is not a power of two. In those cases, at least three steps are required for the binary connection, and the presence of unbalanced leaves increases the cost.
Parallel Baseline methods are consistently slower (up to $1.73\times$), except in one case (1 node to 16 nodes). This overhead stems from spawning extra processes, which in addition leads to oversubscription.
The bottom Figure 4(b) shows shrinking results. Shrinking without spawning processes is clearly advantageous, completing with a speedup of $1387\times$ over the other evaluated approaches. These findings are consistent with previous work [20].
(a) Expansion.
(b) Shrinkage.
Figure 4: Resize times in homogenous system.
Figure 5 summarizes the best method for each ($NS$, NTNT) pair based on statistical tests. Axes are defined with NS on the vertical and NT on the horizontal; thus, the upper triangle corresponds to expansion and the lower triangle to shrinking. When multiple methods appear in a cell, they are statistically equivalent, ordered by ascending time.
The main difference arises in expansions with 8 or fewer groups, requiring at most 3 steps. In such cases, the parallel methods are preferred as they enable shrinking without spawning processes, while maintaining times comparable to Merge, though statistically higher.
Figure 5: Median reconfiguration times in MNV.
5.3 Improved Method Evaluation
On the heterogeneous cluster NASP, only the improved strategy from Section 4.2 was evaluated. Three configurations were tested for expansion and two for shrinking. Each configuration involved a single reconfiguration from $NS$ to $NT$ processes, using up to 72 combinations of nodes from the set 1,2,4,6,8,10,12,14,16{1,2,4,6,8,10,12,14,16}. In all cases, the number of nodes of each type was balanced (half of one type and half of the other). When only one node was used, the 20-core node was selected.
Figure 6 shows the median reconfiguration times as a function of $NS$ and $NT$. The top Figure 6(a) reports expansion results, where the Merge method consistently outperforms the others. The iterative diffusive variant of Merge achieves comparable results, with at most a $1.25\times$ increase, but its cost grows with the number of generated groups.
Baseline methods are always more expensive, ranging from $1.76\times$ to $4.92\times$ slower, due to oversubscription caused by spawning more processes.
The bottom Figure 6(b) shows shrinking results. As in the previous evaluation, Merge is the preferred method with a minimum speedup of $20\times$. To ensure it can be applied, a parallel spawn is required. This approach also benefits future shrink operations, adding only a minor overhead during expansion. However, in difference to the previous evaluation, in 28 out of 32 cases ($87.5\%$), standalone Merge delivered statistically lower times than its iterative diffusive variant.
(a) Expansion.
(b) Shrinkage.
Figure 6: Resize times in heterogenous system.
6 Conclusions and Future Work
This work has introduced and evaluated two new strategies for spawning processes in the traditional MPI. Both approaches rely on parallel spawning with two main goals: (1) reducing the cost of process creation, and (2) enabling efficient shrink operations with the Merge method, the fastest available option. Achieving the second goal requires for each MCW to be fully contained within a node. Previously, this was not possible: when shrinking, processes that should be terminated remained as zombies because other ranks of the same MCW were still active on different nodes, preventing node reuse. The proposed strategies resolve this issue, enabling fast shrink operations.
Between the two strategies, the Hypercube approach is limited to homogeneous, non-shared systems, while the Iterative Diffusive strategy also applies to heterogeneous or shared environments, at the cost of additional computation and memory overhead.
Experimental results on both homogeneous and heterogeneous systems show that, in the worst case, the parallel approaches require up to $1.25\times$ the time of the previous best expansion technique, while enabling shrink operations that are at least $20\times$ faster. In summary, this work provides a practical solution for enabling fast shrink operations without compromising expansion performance.
Future work includes study how to further reduce the synchronization (Subsection 4.3) and connection (Subsection 4.4) overheads, with the aim of outperforming even the basic Merge in expansions. Another promising direction is the design of data redistribution strategies that minimize transfers, ensuring that processes after resizing retain as much of their data as possible.
Declaration of competing interest
The authors have no conflicts of interest to declare that are relevant to the content of this article.
Funding sources
The researchers from UJI have been funded by the project PID2023-146569NB-C22, supported by MCIN/AEI/10.13039/501100011033. Researcher I. Martín-Álvarez was supported by the predoctoral fellowship ACIF/2021/260 from Valencian Region Government and European Social Funds.

### v2

@@
-
+Parallel Spawning Strategies for Dynamic-Aware MPI Applications
 Abstract
 Dynamic resource management is an increasingly important capability of High Performance Computing systems, as it enables jobs to adjust their resource allocation at runtime.
-This capability has been shown to reduce workload makespan, substantially decrease job waiting times and improve overall system utilization.
+This capability can reduce workload makespan, substantially decreasing job waiting times and optimizing resource allocation.
 In this context, malleability refers to the ability of applications to adapt to new resource allocations during execution.
 Although beneficial, malleability incurs significant reconfiguration costs, making the reduction of these costs an important research topic.
-Some existing methods for MPI applications respawn the entire application, which is an expensive solution that avoids the reuse of original processes.
-Other MPI methods reuse them, but fail to fully release unneeded processes when shrinking, since some ranks within the same communicator remain active across nodes, preventing the application from returning those nodes to the system.
-This work overcomes both limitations by proposing a novel parallel spawning strategy, in which all processes cooperate in spawning before redistribution, thereby reducing execution time.
-Additionally, it removes shrinkage limitations, allowing better adaptation of parallel systems to workload and reducing their makespan.
-As a result, it preserves competitive expansion times with at most a $1.25\times$ overhead, while enabling fast shrink operations that reduce their cost by at least $20\times$.
-This strategy has been validated on both homogeneous and heterogeneous systems and can also be applied in shared-resource environments.
+Some existing solutions for MPI applications respawn the entire application, which is an expensive solution that avoids the reuse of original processes.
+Other MPI solutions reuse them, but fail to fully release unneeded processes when shrinking, since some ranks within the same communicator remain active across nodes, preventing the application from returning those nodes to the system.
+This work overcomes both limitations by proposing a novel parallel spawning strategy, in which all processes cooperate in the spawning.
+This allows expansions to reuse processes while also terminating unneeded ones.
+This strategy has been validated on two systems with either machines with equal or different numbers of cores.
+Experiments show that this strategy preserves competitive expansion times with at most a $1.13\times$ and $1.25\times$ overhead for equal and different number of cores per node, respectively.
+More importantly, it enables fast shrink operations that reduce their cost by at least $1387\times$ and $20\times$ in the same scenarios.
 keywords:
+Dynamic Resource Management , Malleability , MPI , Process Management , Heterogenous Allocations
+journal: Journal of Parallel and Distributed Computing
+{graphicalabstract}[Uncaptioned image]
 […]
+[1]organization=Dpto.
+de Ingeniería y Ciencia de los Computadores, Universitat Jaume I, addressline=Av.
+Vicent Sos Baynat, s/n , city=Castelló, postcode=12071, state=Comunitat Valenciana, country=Spain
 1 Introduction
@@
 Scientific applications often underuse their allocations due to factors such as overcommitment, bottlenecks or limited scalability.
-To address this issue, Resource Management Systems (RMS) must be able to distribute resources dynamically, a capability known as Dynamic Resource Management (DRM).
-This capability enables RMSs to adjust job allocations at runtime, provided the applications are malleable, that is, capable of resizing and adapting to changes in resource availability during execution.
+To address this issue, Resource Manager System (RMS) must be able to distribute resources dynamically, a capability known as Dynamic Resource Management (DRM).
+This capability enables RMS s to adjust job allocations at runtime, provided that the applications are malleable, that is, capable of resizing and adapting to changes in resource availability during execution.
 Such flexibility allows applications to achieve more effective optimization of resource utilization [4], throughput [11], energy efficiency [3, 10], and I/O performance [21].
-Recent studies by the US Exascale Computing Project [2] suggest that DRM could significantly enhance the performance of both applications and systems, making it a key contribution for improving relevant metrics in future platforms [13].
+A prior study by the US Exascale Computing Project [2] suggest that DRM could significantly enhance the performance of both applications and systems, making it a key contribution for improving relevant metrics in future platforms [13].
 However, its adoption remains limited compared to other MPI features [6], primarily due to the complexity involved in adapting applications to support malleability and enabling systems to dynamically reallocate resources at runtime.
-There are already libraries available to perform reconfigurations in applications; some of the most relevant are DMR[9], FlexMPI[14], MaM[16] or DPP[8].
+There are already libraries available to perform reconfigurations in applications; some of the most relevant are DPP [8], DMR [11], FlexMPI[16], or MaM [14].
 A particularly challenging scenario for most of these tools is shrinking, i.e., reducing the number of active processes in a job at runtime to free resources or adapt to changing workload demands.
-This operation is limited by the rigid structure of MPI communicators, especially the standard MPI_COMM_WORLD (MCW), which does not allow the partial termination of ranks.
-As shown in [20], terminate-based (TB) methods can shrink without spawning new processes, providing a time efficient alternative to traditional spawn-based (SB) approaches.
-Nevertheless, a major limitation of TB shrinking is the persistence of zombie processes, ranks that remain asleep until the application terminates, as they cannot be fully removed from the global MCW.
+This operation is limited by the rigid structure of the standard MPI communicator MPI_COMM_WORLD (MCW), which does not support the termination of a subset of ranks, even if separated with a MPI_Comm_split function.
+In the paper [17] two methods, Baseline and Merge, are evaluated with different strategies.
+The latter method allowed to shrink the application by the conversion of processes into zombies, which we denote as Zombies Shrinkage (ZS).
+The ZS offered lower reconfiguration times than traditional Spawn Shrinkage (SS) obtained with the Baseline method.
+Nevertheless, a major limitation of ZS is the persistence of zombie processes, ranks that remain asleep until the application terminates, as they cannot be fully removed from the global MCW.
 This creates an issue: the nodes on which these zombies are mapped cannot be returned to the RMS until they terminate, preventing node reuse.
@@
 However, this solution depends on node-specific spawning, which suffers from poor scalability due to its inherently sequential nature [20].
-To overcome these constraints, we propose a novel technique that retains the benefits of TB shrinking while addressing its key limitations, supporting both expansion and shrinkage at minimal cost.
-Our approach isolates each MCW communicator on a separate node and enables all available processes to perform spawning in parallel rather than sequentially, thereby reducing the cost of expansion.
-In addition, it avoids spawning new processes during shrinkage and eliminates the persistence of zombie processes while maintaining the high performance observed in TB methods.
-These improvements make our solution particularly suitable for DRM in large-scale HPC environments, with full support for heterogeneous systems and shared resources.
-The approach has been included in the MaM library [16], which integrates multiple techniques for resizing and data redistribution, allowing the selection of the optimal solution depending on the context.
+Yet, by using this technique it is possible to shrink the application by terminating those processes, an approach that we denote as Termination Shrinkage (TS), which overcomes the key limitation of ZS, its inability to fully terminate processes and make them persist as zombies.
+To overcome the scalability constraints identified in [20], we propose a novel technique that preserves scalability during process spawning while enabling the use of TS for shrinkages.
+Our technique also isolates each MCW communicator on a separate node; however, instead of performing this step sequentially, it is carried out in parallel, allowing each available process to initiate an independent spawn.
+Additionally, newly spawned processes participate in subsequent spawn operations within the same reconfiguration if further spawning is required.
+These improvements make our solution particularly suitable for DRM in large-scale HPC environments, with full support for nodes with different numbers of cores, and for jobs with different numbers of cores on each node.
+Consequently, to evaluate the novel technique, this work considers two types of node allocations: homogeneous, where jobs use the same number of cores in each node, and heterogeneous, where the number of cores used in each node can differ.
+Additionally, the improvements can be used for oversubscription in both types of allocation, where the number of processes exceeds the number of cores allocated to the job.
+Nevertheless, it is not considered the challenge of load balancing between different types of hardware, which is postponed as future work.
+This approach has been included in the MaM library [14], which integrates multiple techniques for resizing and data redistribution, allowing the selection of the optimal solution depending on the context.
 The main contributions of this work are:
-  • The design and implementation of two novel parallel spawning algorithms tailored to the MPI model, enabling scalable malleability for HPC applications.
-  • The extension of the MaM library [16] to support dynamic reconfiguration in heterogeneous queue systems and shared-resource environments.
-  • A comprehensive performance evaluation of the proposed algorithms on both homogeneous and heterogeneous clusters, demonstrating their advantages over the previous best method in terms of scalability and reconfiguration overhead.
+  • The design and implementation of a novel parallel spawning algorithm tailored to the MPI model, enabling scalable malleable executions for HPC applications in homogeneous node allocations.
+  • An adaptation of the previous technique to enable resizing in heterogeneous node allocations.
+  • The capability to shrink malleable applications by terminating processes (TS) without spawning new ones (SS) or leaving them as zombies (ZS).
+  • A comprehensive performance evaluation of the proposed algorithms on jobs with homogeneous node allocations, demonstrating their advantages.
+  • A performance evaluation of the extended algorithm for jobs with heterogeneous node allocations, demonstrating its advantages.
 The remainder of this paper is organized as follows: Section 2 presents a comprehensive review of related work on MPI malleability and spawning techniques.
@@
 Finally, Section 6 concludes the paper and outlines directions for future work.
-2 Background
-According to the classification proposed by Feitelson and Rudolph [5], applications executed in large-scale computing facilities can be grouped into four categories.
-This classification is based on two main criteria: (1) whether the number of processes can change during execution, i.e., whether reconfiguration is supported, and (2) who determines the job size, either the user or the RMS.
-Table 1 summarizes this classification.
+2 Related Work
+According to the classification proposed by Feitelson and Rudolph [5], applications executed in large-scale computing facilities can be grouped into four categories (see Table 1).
+This classification is based on two main criteria: (1) whether the number of processes can change during execution, i.e., whether reconfiguration is supported, and (2) who determines the new job size, either the application or the RMS.
 Job Type Resource Allocation Who Sets Size Description
-Rigid Job Static User Can only be executed with a fixed number of processes defined by the user. No reconfiguration is allowed.
+Rigid Job Static Application Can only be executed with a fixed number of processes defined by the user. No reconfiguration is allowed.
 Moldable Job Static RMS Can start with a variable number of processes. The RMS determines the size when the execution begins.
-Evolving Job Dynamic User Includes a user-defined reconfiguration scheme. The RMS must meet the resource requests for the job to continue executing.
+Evolving Job Dynamic Application Includes a user-defined reconfiguration scheme. The job will only continue if the RMS satisfies the resource request; otherwise, it waits or it is cancelled.
 Malleable Job Dynamic RMS Can be reconfigured at runtime if the RMS decides to adjust the resource allocation.
@@
     Process management.
-The RMS assigns a new number of cores to the job, which dictates how many MPI ranks should be created or terminated.
+The RMS assigns a new number of cores to the job, which determines how many MPI ranks should be created or terminated.
 Processes before resizing are considered sources, while those that continue after resizing are considered targets.
@@
 Incorporating all these stages into a parallel application is a complex task that goes far beyond simple integration with the RMS.
-It requires modifications to the application’s structure to support dynamic creation and termination of processes, manage data redistribution efficiently, and reconfigure communication patterns on the fly.
-These challenges make dynamic applications difficult to implement from scratch, often requiring dedicated tools and frameworks that simplify the transformation of traditional applications into dynamic ones.
+It requires modifications to the application’s structure to support dynamic creation and termination of processes, the efficient management of data redistribution, and the reconfiguration of communication patterns on the fly.
+These challenges make dynamic applications difficult to be implemented from scratch, since they often require dedicated tools and frameworks to simplify the transformation of traditional applications into dynamic ones.
 Without such support, adapting applications to handle resizing becomes both error-prone and time-consuming.
 One malleable solution is Flex-MPI [16], a library built on top of MPICH [19] that bases its spawn operations on the MPI_Comm_spawn function.
-Its primary goal is to help applications become dynamic and the main result of this solution is the Merge method, which is described below.
+The primary goal of this solution is to help applications become dynamic and the main method to resize in Flex-MPI is the Merge method, which is described below.
 When an application expands from $x$ processes to $y$ processes, only the additional yxy-x processes are spawned and added to the existing ones.
@@
 A similar solution is MaM [14], a malleability module also built on top of MPICH and within the MPI scope, developed to analyze and compare different malleability techniques.
-Like Flex-MPI, it bases its spawn operation on the MPI_Comm_spawn function, providing a Merge method, but without the limitation of the minimum number of processes.
-MaM is part of the Proteo framework [19], which helps transform applications into malleable ones, supports the generation of dynamic workloads, and enables performance analysis of different resize operations to select the optimal alternative for a given situation.
+Like Flex-MPI, it bases its spawn operation on the MPI_Comm_spawn function and implements Baseline and Merge methods, in which the former always spawns the new number of processes $y$ and terminates the previous $x$ processes.
+MaM is part of the Proteo framework [18], which helps transform applications into malleable ones.
+It supports the generation of dynamic workloads, and enables performance analysis of different resize operations to select the optimal alternative for a given situation, by simulating different scenarios.
 Another approach is DPP [8], which explores the new concept of MPI Sessions, extending it to implement on-the-fly dynamic mechanisms.
@@
 However, alternative non-standard MPI functions are provided to allow communication between ranks.
-To spawn processes, it uses a custom dynamic RMS, which determines whether a job should get or return resources.
-When a job gets additional resources, it informs a PMIx Reference Runtime Environment (PRRTE) daemon, which then spawns of new processes.
-Once the processes are ready, the job is informed of the new resources and how to connect with them, which is managed by DPP.
-During shrink, processes do not have the limitations of the traditional environment and can be terminated normally.
+DPP uses a custom dynamic RMS, which determines whether a job should get or return resources.
+When a job gets additional resources, it notifies a PMIx Reference RunTime Environment (PRRTE) daemon, which then spawns the new processes.
+Once the processes are ready, the job is notified of the new resources and how to connect with them, which is managed by DPP.
+During shrink, processes do not have the limitations of the traditional environment and can be terminated arbitrarily.
 Lastly, an alternative solution is Dynamic Management of Resources library (DMRlib) [11], which consists of two components: a parallel distributed runtime based on MPI and an extension of Slurm to enable the execution of malleable jobs.
-Initially, the library only allowed the simplest approach, the Baseline method, which always spawns the new number of processes $y$ and terminates the previous $x$ processes.
-However, the library now supports other environments for process management, such as DPP [7] or MaM [10], providing a wider range of possibilities.
-For further details on MPI spawn based malleability, see [1], while for a broader review of dynamic resource management [24] is recommended.
+Initially, the library only allowed the simplest approach, the Baseline method.
+However, the library now supports other environments for process management, such as MaM [9] or DPP [7], providing a wider range of possibilities.
+For further details on MPI spawn based malleability, see [1], while a broader review of dynamic resource management can be found in [22].
 3 MaM overview
@@
 It provides a set of features designed to transform traditional MPI applications into malleable ones, requiring minimal effort from developers.
-Additionally, it defines a simple and intuitive interface, which lowers the barrier to integrating malleability into existing parallel applications [16].
-Specifically, MaM implements multiple methods and strategies to handle the most critical stages of malleability, process management and data redistribution, allowing applications to dynamically adjust their number of MPI ranks during execution.
-For the process management stage, i.e., resizing from an initial group of NS processes to a new group of NT processes, MaM offers two main approaches: the Baseline method, which always creates a new group of processes, and the Merge method, which reuses existing processes and only spawns or terminates the difference in ranks.
+Additionally, it defines a simple and intuitive interface, which makes easier to integrate malleability into existing parallel applications [14].
+Specifically, MaM implements multiple methods and strategies to handle the most critical stages of malleability, process management and data redistribution.
+These techniques allow applications to dynamically adjust the number of MPI ranks during execution.
+Using MaM to incorporate malleability into an application requires specifying a method with none, one, or more of the strategies for process management and data redistribution.
+For the process management stage, i.e., resizing from an initial group of Source processes (NS) to a new group of Target processes (NT), MaM offers two main methods: the Baseline method, which always creates a new group of processes, and the Merge method, which reuses existing processes and only spawns or terminates the difference in ranks.
 These methods can be combined with two strategies: the Asynchronous strategy, which overlaps process creation with application execution to reduce downtime, and the Single strategy, where only one process performs the spawn operation and informs the rest afterwards.
 This flexible design allows users to decide the preferred configuration according to their application needs.
-For the data redistribution stage, MaM distinguishes between constant and variable data, as this affects how and when the redistribution can take place.
-While variable data requires execution to be paused until the transfer is complete, constant data can be redistributed asynchronously, allowing overlap with ongoing computation.
-The library also supports three main communication methods for this stage: Point-to-Point, Collective, or One Sided primitives.
-Additionally, asynchronous redistribution can be implemented through two strategies: Non-blocking communication, which relies on routines like MPI_Test to monitor progress, and Threading, where dedicated threads handle data transfer independently of the main computation.
-Using MaM to incorporate malleability into an application requires specifying a method with none, one, or more of the strategies for stages 22 and $3$.
-A more detailed description of all alternatives related to process management can be found in [20] and to automatic data redistribution in [15] and [18].
-4 Methods for spawning processes in parallel
-As previously discussed, the MaM shrinkage approach is constrained by its inability to fully terminate unneeded processes during execution.
-As a consequence, is not possible to return those resources until the MCW they pertain terminates completely.
-To overcome these issues, this work proposes new strategies based on parallel spawning, which allows processes to be resized efficiently and overcome said situation.
-Parallel process creation is inherently complex and is coordinated across multiple phases: (1) It begins with the spawn of new process groups and opening communication ports.
-(2) it is followed by a synchronization phase, where all groups wait until every required port is ready.
+4 Strategies for spawning processes in parallel
+As previously discussed, the ZS approach is constrained by its inability to fully terminate unnecessary processes during execution.
+As a consequence, it is not possible to return the corresponding nodes until MCW, to which they belong, has terminated completely.
+However, a previous study [17] using a test system showed for shrinkages that the ZS method delivered a speedup of $36\times$ over SS methods.
+Consequently, reaching a solution for shrinkage with TS methods would prove a valuable advantage to reduce the overhead caused by resizing.
+To overcome these challenges, this work proposes new strategies based on parallel spawning, which allows processes to be resized efficiently and create an MCW for each node, facilitating the use of a TS approach.
+Parallel process creation is inherently complex and is coordinated across multiple phases: (1) New processes groups are spawned and communication ports are opened.
+(2) This is followed by a synchronization phase, where all groups wait until every required port is ready.
 (3) Next, the newly spawned groups are merged into a single communicator through the ports.
 (4) Finally, a global reordering of ranks is performed to maintain a logical order between nodes.
-The first phase has been examined through two alternatives: A simpler option only for homogenous systems called Hypercube strategy, and an improved but more complex one that works for heterogenous or shared systems, called Iterative Diffusive strategy.
-The following sections detail each of these steps, highlighting the specific operations involved.
-4.1 Hypercube: Parallel Spawning for Homogeneous Systems
-This parallel strategy is based on performing a call to the MPI_Comm_spawn function for each node that is to be used, assigning all available cores of the node to each newly created process.
+The first phase has been addressed through two alternative approaches: A simpler option only for homogeneous node allocations called Hypercube strategy, and an improved but more complex one that also works for jobs with heterogeneous node allocations, called Iterative Diffusive strategy.
+The following subsections detail these approaches, and explain how the four phases are optimized, highlighting the specific operations involved in each step.
+4.1 Hypercube Strategy
+This parallel strategy is based on invoking the MPI_Comm_spawn function once for each node to be used (up to the total number of required MPI ranks), assigning all available cores of the node (process group) to each newly spawned process.
 The procedure is organized into a series of steps.
-In each step, all active processes, whether source or target, concurrently execute a spawn operation, each one acting on a different node.
-As the number of target processes increases, so does the number of processes participating in the spawning operation.
+In each step, all existing processes concurrently execute a spawn operation, each one acting on a different node.
+As the number of spawned processes increases, so does the number of processes participating in the spawning operation.
 Consequently, with each successive step, a growing number of processes are involved in initiating new spawns, enabling a scalable and distributed expansion across nodes.
@@
 In total, up to $C\cdot I$ processes can be launched, where $C$ is the number of cores per node and $I$ is the number of initial nodes.
-From the second step onward, both the original and previously created processes participate, exponentially increasing the spawning capacity.
+From the second step onward, both the original and previously spawned processes participate, exponentially increasing the spawning capacity.
 However, in the final step, only the first $x$ processes participate, where $x$ corresponds to the remaining number of nodes needed to reach the desired number of nodes.
@@
 Starting with a single fully occupied node, the first step could launch up to $20$ additional nodes.
-In the second step, with 420 processes available (21 nodes with 20 cores each), it would be possible to occupy 410 additional nodes by creating 20 processes per node.
-From a technical standpoint, this strategy is implemented using MPI_Comm_spawn in combination with the MPI_COMM_SELF communicator and an MPI_Info object that specifies the target node for each new group of processes.
-Furthermore, all processes in each node are assigned the group identifier ranging from 0 to the total number of groups to be created (group_id), identifying which group it belongs.
-The number of steps $s$ required to reach a total of $N$ nodes, starting from $I$ initial nodes and with $C$ cores per node, can be calculated using Equation 1.
-$N$ and $I$ are calculated as $\frac{NT}{C}$ and $\frac{NS}{C}$, respectively, where $NT$ is the amount of targets, and $NS$ the amount of sources.
-${s}=\frac{\ln(N/I)}{\ln(1+C)}$ (1)
-From this, the total number of nodes $T_{s}$ at the end of step $s$ can be computed as:
-${T_{s}}=(1+C)^{s}\cdot I$ (2)
-And the corresponding number of processes spawned $t_{s}$ in step $s$ is simply:
-${t_{s}}=T_{s}\cdot C$ (3)
+In the second step, with $420$ processes available ($21$ nodes with $20$ cores each), it would be possible to occupy $420$ additional nodes, creating $20$ processes per node.
+From a technical standpoint, this strategy uses MPI_Comm_spawn in combination with the MPI_COMM_SELF communicator and a MPI_Info object that specifies the target node for each new group of processes.
+Furthermore, all processes in each node are assigned a group identifier ranging from 0 to the total number of groups to be spawned (group_id), identifying which group it belongs.
+From a mathematical perspective, the strategy can be described as a geometric growth, where each existing process expands into $C$ new nodes while the process that creates each node remains active.
+Therefore, the effective growth factor per step is $C+1$.
+The initial number of nodes, $I$, is determined by the total number of source processes NS as:
+$I=\frac{NS}{C}\quad,\quad NS\;mod\;C=0.$
+Using this initial condition and the growth factor, the total number of nodes at step $s$ is given by:
+$T_{s}=\begin{cases}(C+1)^{s}\cdot I-I,&\text{if Method}=\text{Baseline}\\ (C+1)^{s}\cdot I,&\text{if Method}=\text{Merge}\end{cases}\quad,$ (1)
+where $T_{s}$ represents the total nodes at step $s$ and $C+1$ is the growth factor including the cores and the process that spawns them.
+Additionally, it considers the chosen spawning method, either Baseline or Merge.
+For the former, the initial nodes ($I$) are not considered to reach the target nodes ($N$), as the source processes will be removed after the reconfiguration.
+Thus, the corresponding total number of processes at step $s$ is simply obtained by multiplying the number of nodes by the cores per node:
+${t_{s}}=C\cdot T_{s}.$ (2)
+Finally, the total number of steps required to reach a target of $N$ nodes can be derived from Equation 1 by solving for $s$ as:
+$s=\left\lceil\frac{\ln(N/I)}{\ln(C+1)}\right\rceil,$ (3)
+where, $N$ represents the desired total number of nodes, calculated as:
+$N=\frac{NT}{C}\quad,\quad NT\;mod\;C=0,$
+where, NT is the total number of target processes.
 Figure 1 illustrates this strategy with a simple example in which the number of processes is expanded from $NS=1$ to $NT=8$, assuming an architecture with $C=1$ core per node.
-In this case, seven additional groups are created over three steps.
-The process is visualized as a cube, where the edges indicate which group spawns which in each step, and the vertices are labelled with the corresponding group_id, or with the label $I$ for the initial group.
+In this case, seven additional groups are spawned over the course of three steps.
+The process is visualized as a cube, where the edges indicate which group spawns which in each step, and the vertices are labeled with the corresponding group_id, or with the label $I$ for the initial group.
 By the end of the procedure, a total of eight nodes are occupied.
-Figure 1: Generation of 7 groups using parallel spawning. Each step is represented with different markers.
-4.2 Iterative Diffusive: Recursive Algorithm for Heterogeneous Systems
-This method improves upon the previous scheme by including the ability to create process groups of variable sizes at each step, which the Hypercube is not capable of.
-This makes it particularly well-suited to heterogeneous or shared computing environments, where nodes may have different capacities or resources may be distributed unevenly.
-As in the original scheme, the algorithm proceeds in discrete steps, where all existing processes collaborate to spawn new groups on additional nodes.
-The assignment of a unique group identifier (group_idgroup\_id), ranging from 0 to the total number of groups to be created, is also preserved.
-Three vectors are needed to describe the system configuration. The number of entries in these vectors is equal to the number of nodes in the system (N).
-  • $M$: Contains the number of available cores assigned to a job on each node.
-  • AA: Stores the number of processes running on each node before starting this algorithm.
-  • $S$: Has the number of processes to be created in each node when executing this algorithm.
+Figure 1: Generation of $7$ groups using parallel spawning. Each step is represented with different markers, and the number marks the group identifier.
+This strategy removes the shared MCW when expanding across multiple nodes. Therefore, it enables shrinkage via the TS method by releasing the excess processes.
+4.2 Iterative Diffusive Strategy
+The main limitation of the Hypercube strategy is that all spawned process groups have a constant size.
+The Iterative Diffusive parallel strategy can use variable sizes for each process group.
+This makes it particularly well-suited for jobs with heterogeneous node allocations.
+As in the previous scheme, the algorithm proceeds in discrete steps, where all existing processes collaborate to spawn new groups on additional nodes.
+The assignment of a unique group identifier (group_idgroup\_id), ranging from 0 to the total number of groups to be spawned, is also preserved.
+Three vectors are needed to describe the allocation configuration, and are maintained through the execution.
+The number of entries in these vectors is equal to the number of nodes in the allocation ($N$).
+  • AA: Contains the total number of cores assigned to the job on each node.
+  • $R$: Stores the number of processes of the job running on each node.
+  • $S$: Keeps the number of processes of the job to be spawned in each node when executing this algorithm. This vector is computed from the previous ones as follows:
+$S_{i}=A_{i}-R_{i},\quad\text{for }i=0,1,\dots,N-1.$
 Based on these vectors, a stepwise sequence can be constructed to determine the total number of processes and nodes involved at each spawning step.
-Only nodes with positive entries in the vector $S$ will allocate new processes; thus, each non-zero entry indicates the creation of a new group on the corresponding node.
-This vector is computed from the previous ones as follows:
-$S_{i}=M_{i}-A_{i},\quad\text{for }i=0,1,\dots,N-1$
-Equation 4 defines the total number of processes $t_{s}$ present in step $s$, as a function of the cumulative total of previous steps ($t_{s-1}$) and the number of processes generated in the current step ($g_{s}$).
-Step $s=0$ corresponds to the initial state, which includes only the source processes, as specified by the vector $A$.
-$t_{s}=\begin{cases}\sum\limits_{j=0}^{N-1}A_{j},&\text{if }s=0\\ t_{s-1}+g_{s},&\text{if }s>0\end{cases}$ (4)
-Equation 5 defines the number of processes $g_{s}$ generated in step $s$.
-At each stage, the elements of vector $S$ from index $\nu_{s}$ to index $\lambda_{s}$ are accumulated to calculate the number of new processes that need to be generated.
-$g_{s}$ $=\sum\limits_{j=\nu_{s}}^{\lambda_{s}}S_{j}$ $\quad\text{where}\quad\lambda_{s}$ $=\min\left(N-1,\,\left(\sum\limits_{i=0}^{s-1}t_{i}\right)-1\right),$ (5)
-$\nu_{s}$ $=\sum\limits_{i=0}^{s-2}t_{i}$
-Finally, Equation 7 represents the cumulative number of occupied nodes ($T_{s}$) up to step $s$, whereas Equation 6 specifies the number of new nodes ($G_{s}$) occupied by newly created groups during that step.
+Note that only nodes with positive entries in the vector $S$ will allocate new processes; thus, each entry greater than zero indicates the creation of a new group on the corresponding node.
+The number of existing processes in each step $s$ can then be calculated as follows:
+$t_{s}=\begin{cases}\sum\limits_{j=0}^{N-1}R_{j},&\text{if }s=0\\ t_{s-1}+g_{s},&\text{if }s>0\end{cases}\quad,$ (4)
+where $t_{s}$ is the total processes existing at the end of step $s$, calculated as a function of the cumulative total of previous steps ($t_{s-1}$) and the number of processes generated in the current step ($g_{s}$).
+The initial state ($s=0$) corresponds to the number of source processes specified in vector $R$, and the number of generated processes ($g_{s}$) in a given step can be calculated as the sum of a sequence of elements in vector $S$:
+$g_{s}=\sum\limits_{i=\lambda_{s-1}}^{\min\left(N,\,\lambda_{s}\right)-1}S_{i},$ (5)
+where the limits $\lambda_{s}$ defines the range of indices of vector $S$ used at each spawning step, as follows:
+$\lambda_{s}=\begin{cases}0,&\text{if }s=0\\ \lambda_{s-1}+t_{s-1},&\text{if }s>0\end{cases}\quad.$ (6)
+These limits ensure that each step consumes a contiguous segment of $S$, without overlap or omission, while maintaining a sequential progression through the vector.
+Specifically, $\lambda_{s-1}$ marks the starting index for step $s$, which corresponds to the total number of processes handled in all previous steps.
+Conversely, $\lambda_{s}-1$ marks the last index to be included in the current step, extending over the $t_{s-1}$ processes that participate at this stage.
+As a result, each step $s$ uses exactly $t_{s-1}$ entries of $S$, starting right after the portion used in step s1s-1.
+The minimum in the upper index for Equation 5 is used to ensure no more than $N$ spawns are performed.
 In all equations, null elements in vector $S$ are disregarded.
-$G_{s}=\min(N-T_{s-1},\,t_{s-1})$ (6)
-$T_{s}=\begin{cases}I,&\text{if }s=0\\ T_{s-1}+G_{s},&\text{if }s>0\end{cases}$ (7)
+The cumulative number of occupied nodes ($T_{s}$) up to step $s$ is calculated from these previous equations as follows:
+$T_{s}=\begin{cases}I,&\text{if }s=0\\ T_{s-1}+G_{s},&\text{if }s>0\end{cases}\quad,$ (7)
+where for each step, the sum of the nodes from the previous step ($T_{s-1}$) and the nodes added during the current step ($G_{s}$) are calculated.
+In the first step ($s=0$) the initial number of nodes is used ($I$).
+For the next steps the number of nodes mainly depends on $G_{s}$, which represents the number of new nodes to add.
+A new node is added if the corresponding positions in the vectors $R$ and $S$ do not contain any processes before spawning ($R_{i}=0$) and if there are processes to spawn ($S_{i}>0$).
+Therefore, one node is added for each position considered in the range defined by Equation 6, only if both conditions hold.
+This can be represented as follows:
+$G_{s}=\sum\limits_{i=\lambda_{s-1}}^{\min\left(N,\,\lambda_{s}\right)-1}\{R_{i}=0\wedge S_{i}>0\},$ (8)
+where the limits $\lambda_{s}$ are calculated by using Equation 6. When the condition of the summation is met, it adds 1, or 0 otherwise.
+Table 2 shows how the different Equations behave depending on the initial parameters $A$, $R$, $S$, NN and $I$. In this example, $s=0$ is the initial state before executing this strategy.
+Initial Vectors
+A [4, 2, 8, 12, 3, 3, 4, 4, 6, 3]
+R [2, 0, 0,  0,  0, 0, 0, 0, 0, 0]
+S [2, 2, 8, 12, 3, 3, 4, 4, 6, 3]
+Computed Values per Step
+$s$ $t_{s}$ $g_{s}$ $\lambda_{s}$ $T_{s}$ $G_{s}$
+0 2 - 0 1 -
+1 6 4 2 2 1
+2 40 34 7 8 6
+3 49 9 47 10 2
+Table 2: Iterative diffusive procedure example for a new allocation of resources from 11 node (II) to $10$ nodes ($N$).
+Similar to the previous approach, this one ensures that no shared MCW exists across the expanded nodes, with the added advantage of accounting for a heterogeneous number of cores on each expanded node.
+Moreover, it also enables fast shrinkage through TS method.
 4.3 Synchronization Between Process Groups
-Once all processes have been created, it is necessary to establish connections between them to enable direct communication.
-With the adopted creation strategy, each newly spawned group initially knows only its parent process (the one that created it) and, if applicable, its own child processes.
-As shown in Figure 1, processes can only communicate directly with those to which they are connected by an edge in the spawn graph.
+The purpose of synchronizing the groups is to ensure that all processes know that every port is active before any connection attempt is made.
+Once all processes have been spawned, establishing these connections is essential, as they enable direct communication among them.
+Due to the chosen process creation strategies, each newly spawned group only knows its parent process (its creator) and, if applicable, its own child processes.
+Consequently, as shown in Figure 1, processes can only communicate directly with those linked to them by an edge in the spawn graph.
 To enable communication between separate groups, the functions MPI_Comm_connect and MPI_Comm_accept are used.
@@
 First, the process groups must be prepared for connection.
-The processes designated as roots in the MPI_Comm_accept calls must first open a port using MPI_Open_port and publish the service name via MPI_Publish_name.
+Thus, the processes designated as roots in the MPI_Comm_accept calls must first open a port using MPI_Open_port and publish the service name via MPI_Publish_name.
 Since MPI_Comm_connect must be called after the corresponding port has been opened by the other group, a global synchronization phase between all groups is required.
-Otherwise, execution errors may occur, as observed with implementations such as MPICH [^1] This behaviour has only been tested with MPICH..
-The synchronization between groups consists of three steps and is carried out within each group using a dedicated subcommunicator.
-The objective is to ensure that all processes are aware of the port status before attempting to establish a connection.
-  1. [1]
-  2. 1.
+Otherwise, execution errors may occur, as observed with implementations such as MPICH [^1] This behavior has only been tested with MPICH..
+For the rest of the paper, the root process of each group is always the process with rank 0 in its communicator, which is required by the functions discussed here.
+Each group carries out the synchronization using a dedicated subcommunicator consisting of three stages.
+  1. 1
     Subcommunicator creation.
-Each group creates a subcommunicator using MPI_Comm_split, including the root process of the group and all processes that have created child groups.
+Each group creates a subcommunicator using MPI_Comm_split, including the root process of the group and all processes of the group that have spawned child groups.
 This subcommunicator will be used to coordinate intra-group synchronization.
-  3. 2.
+  2. 2
     Upside Synchronization.
-Each process that has spawned child groups waits to receive a token from each of them using MPI_Irecv followed by MPI_Waitall.
+Each process that has spawned child groups waits to receive a token from each of them using MPI_Irecv and MPI_Waitall.
 Then, all processes in the subcommunicator execute a call to MPI_Barrier, ensuring that all tokens have been received.
 Finally, the root process (if not part of the source group) notifies its parent group using MPI_Send.
-  4. 3.
-    Downside Synchronization The root process of each group (except the source group) waits to receive a token from its creator via MPI_Recv.
+  3. 3
+    Downside Synchronization.
+The root process of each group (except the source group) waits to receive a token from its creator via MPI_Recv.
 Then, processes in the subcommunicator perform a MPI_Barrier call (excluding the source group).
-Finally, each process that has spawned child groups sends them a token using MPI_Isend followed by MPI_Waitall.
-These barriers are essential to ensure that no group proceeds before all of its descendants have completed the Upside step, and to prevent any group from notifying its children before being notified by its parent group.
-Listing 1 shows the implementation of the synchronization mechanism.
-The processes use the world_c communicator, which corresponds to MCW for spawned processes, or to a shared communicator among source processes in the case of creators.
-The parent_c communicator is an intercommunicator between a spawned group and its parent.
-The variable qty_s indicates the number of child groups spawned by a process, and the corresponding intercommunicators are stored in the spawn_c vector.
+Finally, each process that has spawned child groups sends them a token using MPI_Isend, followed by MPI_Waitall.
+The barriers included in the stages are essential to ensure that no group proceeds before all its descendants have completed the Upside step, and to prevent any group from notifying its children before being notified by its parent group.
+However, the Waitall is required only to ensure the communication is completed.
+The left and right sides of Figure 2 illustrate, respectively, the upside synchronization and downside synchronization steps.
+In this example, six process groups are spawned in two spawning steps.
+The arrows indicate which processes send messages to signal their readiness during the upside phase, and to confirm that all groups are ready in the downside phase.
+A group represented in black with stripes indicates that it must wait to receive all messages before proceeding.
+At that point, a MPI_Barrier is used to synchronize the group, after which the execution can continue.
+Listing 1 shows the implementation of the synchronization mechanism, where the first stage is shown in lines L13-L17, the second from lines L19-L28, and the third from lines L30-L41.
+The processes use the world_c communicator, which corresponds to MCW for spawned processes or to the global communicator among source processes.
+Then, the parent_c communicator is an inter-communicator between a spawned group and its parent.
+Additionally, the variable qty_c indicates the number of child groups spawned by the calling process, and the corresponding inter-communicators are stored in the spawn_c vector.
+A particular MPI_Barrier function is needed at line L24, because other ranks of the same group could also have children, and therefore the group has to ensure all messages have been received before continuing.
+This is represented for the Init Node group and the Group_Id 0 group in the upside synchronization in Figure 2.
 1void common_synch(MPI_Comm world_c, MPI_Comm parent_c,
-2 int qty_s, MPI_Comm *spawn_c) {
-3 int myId, i, root = 0;
-4 int color = qty_s ? 1 : MPI_UNDEFINED;
-5 MPI_Request *reqs = malloc(qty_s * sizeof *reqs);
+2 int qty_c, MPI_Comm *spawn_c) {
+3 // world_c: Built comm for sources, MCW for targets
+4 // parent_c: Intercomm to contact parent rank if any
+5 // qty_c: Total children groups for calling rank
+6 // spawn_c: Children comms array for calling rank
+7 int rank, i, root = 0;
+8 // The color separates ranks into two subgroups
+9 int color = qty_c ? 1 : MPI_UNDEFINED;
+10 MPI_Request *reqs = malloc(qty_c * sizeof(MPI_Request));
 11 MPI_Comm synch_ranks;
-7
+12
+13// Stage 1: Subcommunicator creation
 14 // Only root ranks or ranks with children synchronize
-9 MPI_Comm_rank(world_c, &myId);
-10 MPI_Comm_split(world_c, color, myId, &synch_ranks);
-11
+15 // synch_ranks is only used by ranks with children
+16 MPI_Comm_rank(world_c, &rank);
+17 MPI_Comm_split(world_c, color, rank, &synch_ranks);
+18
+19// Stage 2: Upside Synchronization
 20 // Wait for all children of this group
-13 for(i=0; i<qty_s; i++) {
-14 MPI_Irecv(…, source=root, spawn_c[i], &reqs[i]);
+21 for(i=0; i<qty_c; i++) {
+22 MPI_Irecv(…, root, spawn_c[i], &reqs[i]);
 23 }
-16 if(qty_s){MPI_Waitall(reqs);MPI_Barrier(synch_ranks);}
-17
-18 if(parent_c != MPI_COMM_NULL && myId == root) {
-19 // Send to parent this group is ready
-20 MPI_Send(…, dest=root, parent_c);
+24 if(qty_c){ MPI_Waitall(reqs); MPI_Barrier(synch_ranks); }
+25 if(parent_c != MPI_COMM_NULL && rank == root) {
+26 // Send to parent that this group is ready
+27 MPI_Send(…, root, parent_c);
+28 }
+29
+30// Stage 3: Downside Synchronization
+31 if(parent_c != MPI_COMM_NULL && rank == root) {
 32 // Wait for the other groups readiness
-22 MPI_Recv(…, source=root, parent_c);
+33 MPI_Recv(…, root, parent_c);
 34 }
-24
-25 // Notify this group that all groups are ready
-26 if(parent_c != MPI_COMM_NULL && qty_s)
+35 // Notify all ranks in the group to continue
+36 if(parent_c != MPI_COMM_NULL && qty_c)
 37 {MPI_Barrier(synch_ranks);}
 38 // Notify children that all groups are ready
-29 for(i=0; i<qty_s; i++)
+39 for(i=0; i<qty_c; i++)
 40 {MPI_Isend(…, dest=root, spawn_c[i], &reqs[i]);}
-31 if(qty_s) {MPI_Waitall(reqs);}
-32
+41 if(qty_c) {MPI_Waitall(reqs);}
+42
 43 if(synch_ranks != MPI_COMM_NULL)
 44 {MPI_Comm_disconnect(&synch_ranks);}
-35}
-Listing 1: Syncrhonization step before binary connection.
-Figure 2 illustrates the upside synchronization and downside synchronization steps, shown on the left and right sides, respectively.
-In this example, six process groups are created in two spawning steps.
-The arrows indicate which processes send messages to signal their readiness in the upside phase, and to confirm that all groups are ready in the downside phase.
-A group represented in black with stripes indicates that it must wait to receive all messages before proceeding.
-At that point, a MPI_Barrier is used to synchronize the group, after which the execution can continue.
-Figure 2: Synchronization of 6 spawned groups and the initial node. The arrows indicate the direction of messages, and stripes the need of that group of using a Barrier.
+45}
+Listing 1: ]Synchronization step before binary connection.
+Figure 2: Synchronization of $6$ spawned groups and the initial node. The arrows indicate the direction of messages, and the stripes the necessity of using Barriers.
 4.4 Binary connection
-Once synchronization between all groups is complete, the connection operation can begin.
-The adopted strategy consists of making pairwise binary connections in several successive steps.
-In each step, groups with an identifier less than half the number of active groups perform an MPI_Comm_accept operation, while those with a higher identifier execute an MPI_Comm_connect.
+Once synchronization between all groups is completed, the connection operation can begin.
+The adopted strategy consists of establishing binary connections between pairs of groups in several successive steps.
+In each one, groups with an identifier less than half the number of active groups perform an MPI_Comm_accept, while those with a higher identifier execute an MPI_Comm_connect.
 In the case of an odd number of groups, the middle group is temporarily excluded from that step and proceeds directly to the next one.
-Each group connects to the group whose identifier is its mirror image within the current set of active groups, that is, the group symmetrically positioned relative to the center.
 Once the connection is completed, both groups are merged into a new communicator and adopt a common group identifier, corresponding to the group that executed the MPI_Comm_accept call.
@@
 Groups performing accept simply wait for incoming connections, regardless of the order in which they arrive.
-Figure 3: Connection of 7 spawned groups in 3 steps. In each step the amount of groups is halved and connections operations may be unordered.
+Figure 3: Connection of $7$ spawned groups in 33 steps. In each step the number of groups is halved and connections operations may be unordered.
 Listing 2 presents a simplified version of the code used to implement this binary connection strategy between groups.
-In the code, the groups are distinguished based on whether they should perform a MPI_Comm_accept operation (lines L11–L15) or a MPI_Comm_connect (lines L17–L23).
-In the latter case, the group obtains the corresponding port name (line L19) using the group_id of the target group and the MPI_Lookup_name function.
+In the code, the groups are distinguished based on whether they should perform an MPI_Comm_accept operation (lines L15–L19) or an MPI_Comm_connect (lines L20–L26).
+In the latter case, the group obtains the corresponding port name by calling the get_remote_port function (line L22), which uses the target group’s group_id to return the port name.
+To do so, it first writes an established service name based on the group_id, and then calls MPI_Lookup_name, which retrieves the port name associated with the service.
 1void binary_connection(int groups, int group_id,
 2 char *my_port, MPI_Comm *new_comm) {
-3 int middle, new_groups, new_gid;
+3 // groups: Total spawned groups.
+4 // group_id: Group identifier for calling rank
+5 // my_port: Local port name of calling rank if any
+6 // new_comm: Output communicator of combined groups
+7 int middle, new_groups, new_gid, root = 0;
 8 char *port;
@@
 13 middle = groups / 2;
-10 new_groups = ceil(groups / 2.0);
+14 new_groups = groups - middle;
 15 if(group_id < middle) {
-12 MPI_Comm_accept(my_port, MPI_INFO_NULL,
-13 MAM_ROOT, merge_comm, &intercomm);
+16 MPI_Comm_accept(my_port, MPI_INFO_NULL, root,
+17 merge_comm, &intercomm);
 18 MPI_Intercomm_merge(intercomm, 0, &aux_comm);
 19 merge_comm = aux_comm;
-16
 20 } else if(group_id >= new_groups) {
@@
 22 get_remote_port(new_gid, &port);
-20 MPI_Comm_connect(port, MPI_INFO_NULL,
-21 MAM_ROOT, merge_comm, &intercomm);
+23 MPI_Comm_connect(port, MPI_INFO_NULL, root,
+24 merge_comm, &intercomm);
 25 MPI_Intercomm_merge(intercomm, 1, &aux_comm);
 26 merge_comm = aux_comm; group_id = new_gid;
-24
+27 }
 28 groups = new_groups;
+29 }
 30 *new_comm = merge_comm;
-27}
-Listing 2: Binnary connection step for the parallel spawn.
-4.5 Rank Reordering and Connection with sources
-Once all the generated groups have been merged into a single one, the process ranks may not be ordered across nodes.
-This is because binary connections between groups do not enforce a specific order and are susceptible to race conditions.
-To ensure a consistent global order, an MPI_Comm_split call is made, where each process specifies its new rank as follows:
-$\text{world\_rank}+\sum\limits_{j=0}^{N-1}A_{j}+\sum\limits_{j=0}^{\text{group\_id-1}}S_{j}$
-At the same time, all intermediate communicators created throughout the strategy are disconnected using MPI_Comm_disconnect.
-This frees up resources and ensures that any node can be removed or replaced in future reconfigurations without leaving residual dependencies.
-Finally, the target group is connected to the source group, establishing the final inter-communicator that links them. This operation concludes the process creation and connection phases.
+31}
+Listing 2: ]Binary connection step for the parallel spawn.
+4.5 Rank Reordering
+Once all the generated groups have been merged into a single one, the MPI ranks may not be ordered across nodes.
+This is because binary connections between groups do not enforce a specific order and, therefore, are susceptible to race conditions.
+To ensure a consistent global order, an MPI_Comm_split call is made in which each process specifies its new rank as follows:
+$\text{world\_rank}+\sum\limits_{j=0}^{N-1}R_{j}+\sum\limits_{j=0}^{\text{group\_id-1}}S_{j},$ (9)
+where, world_rank is the rank that the calling process had in its MCW, while both summations take into account all the previous ranks.
+The first one counts the number of ranks that existed before the resize, but is not computed since the value is obtained when the group is spawned and can be used directly.
+Meanwhile, the second summation counts the number of ranks in all groups whose group_id is smaller than that of the calling rank.
+4.6 Overview
+To conclude this subsection, a global overview of the tasks required to perform the parallel spawn is provided, both for source and newly spawned processes.
+The steps for sources must be carried out in the following order:
+  1. 1.
+    Open a port and service name for the root rank.
+  2. 2.
+    Use one of the strategies described in Subsections 4.1 and 4.2 to perform the parallel spawn.
+  3. 3.
+    Synchronize ranks to ensure correctness during the binary connection. (Subsection 4.3)
+  4. 4.
+    Connect the merged-groups communicator with the initial group communicator.
+Listing 3 shows these tasks for the source processes.
+First, they compute the required initial data (line L14), and then the root process opens a port (line L17).
+At this point, the parallel spawn begins, selecting the appropriate strategy depending on whether the new node allocation is homogeneous or heterogeneous (lines L20–L24).
+Note that, each algorithm determines whether a given rank must spawn additional groups.
+Next, the ranks synchronize to ensure correctness (line L27).
+Afterwards, unnecessary communicators are disconnected (lines L30-L31), and finally the initial group waits for the merged group of spawned processes to connect to the port opened by the root process (lines L32-L34).
+In the case of newly spawned processes, the steps to carry out are the following:
+  1. 1.
+    Open a port for root ranks whose group identifier satisfies that group_id<total_spawned_groups/2{group\_id}<{total\_spawned\_groups}/2. (Prior preparation to Subsection 4.4)
+  2. 2.
+    Use one of the strategies described in Subsections 4.1 and 4.2 to perform the parallel spawn.
+  3. 3.
+    Synchronize ranks to ensure correctness during the binary connection. (Subsection 4.3)
+  4. 4.
+    Carry out the binary connection to merge the spawned groups into a single one. (Subsection 4.4)
+  5. 5.
+    Reorder the ranks to match the expected ordering. (Subsection 4.5)
+  6. 6.
+    Connect the merged-groups communicator with the initial group communicator.
+Listing 4 shows the tasks carried out for the generated groups.
+At startup, they first obtain the data they will use, which they receive from their parent group upon creation (lines L17-L18).
+Then, in each group, the root rank of those groups whose group_id is lower than half of the generated groups (groups–I) proceeds to open a port (lines L21–L22).
+At this point, the parallel spawn begins, selecting the appropriate strategy depending on whether the new node allocation is homogeneous or heterogeneous (lines L25–L29).
+Once all groups have been created, they synchronize to ensure that all ports have been opened (line L32), and then disconnect any communicators that are no longer needed (lines L33–L36).
+After this synchronization, all spawned processes are merged into a single communicator (lines L39-L40) and then reordered according to the expected rank ordering (line L41).
+Finally, the root process of the merged group obtains the port name of the initial group (lines L44-L45) and connects to it (lines L46-L47).
+1void parallel_strat_parents(MPI_Comm group_comm, int rank,
+2 MPI_Comm *out_comm) {
+3 // group_comm: Built comm for sources
+4 // rank: Identifier of calling process
+5 // out_comm: New intercomm for sources and spawned
+6 int qty_c;
+7 int groups, root;
+8 int N, *A, *R, *S;
+9 char *port_name;
+10 MPI_Comm *spawn_c = NULL;
+11
+12 qty_c = 0; root = 0;
+13 // Calculate following data
+14 get_initial_data(&N,&A,&R,&S, &groups);
+15
+16 //Open port if rank 0 (root)
+17 open_port(&port_name, !rank, groups);
+18
+19 // Decide spawn depending on core counts
+20 if(check_homogenous_dist(S)) {
+21 hypercube(…, &qty_c, &spawn_c);
+22 } else {
+23 diffusive_iterative(…, &qty_c, &spawn_c);
+24 }
+25
+26 // Synchronize all ranks for the binary connection
+27 common_synch(group_comm, MPI_COMM_NULL, qty_c, spawn_c);
+28
+29 // Create intercomm between sources and children
+30 for(int i=0; i<qty_c; i++)
+31 {MPI_Comm_disconnect(&spawn_c[i]);}
+32 if(spawn_c != NULL) free(spawn_c);
+33 MPI_Comm_accept(port_name, MPI_INFO_NULL, root,
+34 group_comm, out_comm);
+35}
+Listing 3: ]Overall tasks for source ranks.
+1void parallel_strat_children(int rank, MPI_Comm *out_comm,
+2 MPI_Comm *newintracomm) {
+3 // rank: Identifier of calling process
+4 // out_comm: New intercomm for sources and spawned
+5 // newintracomm: Output intracomm for spawned ranks
+6 int i, exp_id, group_id, opening, qty_c;
+7 int groups, root;
+8 int N, I, *A, *R, *S;
+9 int initial_group_id = -1;
+10 char *port_name, *remote_port_name;
+11 MPI_Comm *spawn_c = NULL;
+12 MPI_Comm group_comm = MPI_COMM_WORLD;
+13 MPI_Comm_get_parent(out_comm);
+14
+15 qty_c = 0; root = 0;
+16 // Data is obtained when rank is spawned
+17 get_initial_data_children(&N,&A,&R,&S, &groups, &I,
+18 &group_id);
+19
+20 // Open port if required for binary connection
+21 opening = (rank == root && group_id < (groups-I)/2) ?1:0;
+22 open_port(&port_name, opening, group_id);
+23
+24 // Decide spawn depending on core counts
+25 if(check_homogenous_dist(S)) {
+26 hypercube_spawn(…, &qty_c, &spawn_c);
+27 } else {
+28 diffusive_iterative_spawn(…, &qty_c, &spawn_c);
+29 }
+30
+31 // Synchronize all ranks for the binary connection
+32 common_synch(group_comm, *out_comm, qty_c, spawn_c);
+33 for(i=0; i<qty_c; i++)
+34 { MPI_Comm_disconnect(&spawn_c[i]); }
+35 if(spawn_c != NULL) free(spawn_c);
+36 MPI_Comm_disconnect(out_comm);
+37
+38 // Connect groups and ensure expected rank order
+39 binary_connection(groups - I, group_id, port_name,
+40 newintracomm);
+41 rank_reorder(newintracomm, N, R, S, rank);
+42
+43 // Create intercomm between sources and children
+44 if(rank == root)
+45 get_remote_port(initial_group_id, &remote_port_name);
+46 MPI_Comm_connect(remote_port_name, MPI_INFO_NULL, root,
+47 *newintracomm, out_comm);
+48}
+Listing 4: ]Overall tasks for target ranks.
+In cases of oversubscription, where the number of active processes exceeds the number of cores allocated to the job, both of the proposed strategies can be applied.
+In this context, the vector $A$ can be adjusted to reflect the expected level of oversubscription on each node.
+However, for the first strategy, it is necessary to ensure that all non-zero entries of $A$ are equal across nodes.
+An additional important consideration appears for the initial process group that starts the application.
+When the execution begins using more than one node, the resulting MCW cannot returns nodes partially.
+Within the MaM library, there are two possible approaches to handle this situation:
+  • Detect this condition during MaM initialization and recreate all processes using the Baseline method together with the parallel spawning strategy described in this work; or
+  • Detect the issue and postpone any corrective action until it is actually required.
+In this work, the second approach has been adopted.
+The reason is that the application may never need to reduce its resource usage.
+If a shrink is requested, the action to be taken depends on the execution history:
+  • No previous expansions have been performed.
+The application still relies entirely on the initial MCW, which cannot be partially returned.
+Therefore, in case a future shrink is required, MaM performs a parallel expansion to enable TS.
+  • At least one expansion has been performed. Two cases may arise depending on the requested reduction size:
+    • The number of nodes to be returned is smaller than the original allocation.
+Only the nodes allocated in previous expansions are returned, preserving the initial MCW intact.
+The issue is postponed.
+    • The number of nodes to be returned is greater than or equal to the original allocation.
+The entire initial MCW can be safely released.
+If additional nodes must be returned, they come from the expanded set as well.
+In this scenario, the challenging MCW disappears entirely.
+  • The RMS requests the release of a subset of cores within a node. The excess ranks can be turned into zombie processes, enabling a partial shrink with ZS [17].
+4.7 Enabling a TS shrink
+To ensure a TS can be used instead of ZS, some information must be recorded in the root rank of the global communicator and in the root rank of each MCW.
+The root rank of the global communicator must maintain a structure that contains for each MCW, the nodelist where they are executing.
+This structure enables the root to decide which approach of Subsection 4.6 should be used when shrinking and decide which groups to terminate.
+Additionally, if the root rank decides that its own group should be terminated: first, it selects as the new root rank the rank whose identifier is the lowest between the ones that will stay after the reconfiguration; and second it must communicate the structure to that new root rank of the global communicator.
+At the MCW level, the root rank maintains a structure that records, for each rank in the communicator, whether it is active or marked as a zombie, as well as the node on which it resides.
+When a partial shrink affecting only a subset of cores within a node is requested, the corresponding ranks are marked as zombies and enter a sleep state.
+If, instead, all ranks in the MCW are marked as zombies, the group transitions to a TS operation: the zombie ranks are awakened and all processes in the communicator terminate.
+If the RMS requests the release of all nodes associated with an MCW, the group is terminated directly using TS.
+However, if an MCW spans multiple nodes and the RMS requests the release of only a subset of those nodes, TS cannot be applied.
+In this case, ranks cannot be selectively marked as zombies, and the root of the global communicator must fall back to an alternative shrink strategy based on ZS.
 5 Experimental results
+The experiments have been designed to address the following questions:
+  • Do the proposed parallel strategies reduce the resize time when expanding in comparison with the Merge method?
+  • How do these strategies behave in terms of resize time during shrink operations when compared to a TS method?
+The first question evaluates whether the proposed strategies provide an improvement over the Merge method.
+Even if no improvement is observed, it is still necessary to assess whether the additional overhead is justified, given that the primary goal of the parallel strategies is to enable the use of the extremely fast TS method for shrinkage.
+The second question addresses two objectives.
+Firstly, it quantifies the difference in resize time between the parallel strategies and the TS method, highlighting the effectiveness of the latter.
+Secondly, it enables direct comparison with results previously reported in the literature [17].
 5.1 Hardware and Software
 The experiments were conducted on two clusters.
-The first one, MarenostrumV [^2] https://www.bsc.es/supportkc/docs/MareNostrum5/overview (MNV), used 32 nodes from the general queue.
+The first one, Marenostrum 5 (MN5)  [^2] https://www.bsc.es/supportkc/docs/MareNostrum5/overview, used $32$ nodes from the general queue.
 Each node has two 56-core Intel Xeon 8480 CPU s ($3584$ cores in total).
-This system was employed to analyze the behaviour of both spawning techniques described in Section 4.
+This system was employed to analyze the behavior of both spawning strategies described in Section 4.
 The second cluster, Nasp [^3] https://www.hpca.uji.es/computing-facilities/, consists of two sets of 8 nodes.
-The first set has nodes with two 10-core Intel Xeon 4210 CPUs (160 cores total), connected through a 100 Gbps InfiniBand EDR and a 10 Gbps Ethernet network.
-The second set uses nodes with 32-core Intel Xeon 6346 CPUs (256 cores total) connected via 10 Gbps Ethernet.
-Both sets communicate through a shared 10 Gbps Ethernet link between their switches.
-This heterogenous cluster was used to evaluate the effectiveness of the improved method in Section 4.2.
-All experiments in MNV were run with MPICH 4.2.0 and CH4:OFI (InfiniBand), while in NASP with MPICH 3.4.3 and CH3:Nemesis (Ethernet).
+The first set has nodes with two 10-core Intel Xeon 4210 CPU s ($160$ cores total), connected through a 100 Gbit/s InfiniBand Endpoint Detection and Response (EDR) and a 10 Gbit/s Ethernet network.
+The second set uses nodes with 32-core Intel Xeon 6346 CPU s ($256$ cores total) connected via 10 Gbit/s Ethernet.
+Both sets communicate through a shared 10 Gbit/s Ethernet link between their switches.
+This heterogeneous cluster was used to evaluate the effectiveness of the heterogeneous-aware strategy in Section 4.2.
+All MN5 experiments were run with MPICH 4.2.0 and CH4:OFI (InfiniBand), while NASP experiments were run with MPICH 3.4.3 and CH3:Nemesis (Ethernet).
+Proteo was used to evaluate the different approaches. Before reconfiguration, 5 iterations of Monte Carlo Pi computation including one MPI_Allgather were performed to ensure MPI initialization.
+For expansions, MaM was configured using Baseline (B) and Merge (M) methods, each combined with the two strategies described in Section 4.
+All these configurations are compared against Merge without strategies, which was shown to be the best expansion method in a previous study [17].
+For shrinking, MaM was configured to execute B combined with both strategies described in Section 4, and these configurations were compared against the Merge method without strategies.
+When shrinking, the Merge method in MaM corresponds now to the TS method [^4] Previously, it was the ZS method, as it was not possible to fully terminate zombies.; therefore, since no processes are spawned, it cannot use the described parallel strategies.
+Nevertheless, thanks to the use of parallel strategies in previous resizes, it ensures that TS will return the nodes to the RMS instead of creating zombies.
+Moreover, since the ZS method provided the best resizing times in a previous work [17], its counterpart TS is used here both to complement the results of that study and to show the benefits of TS.
+Each configuration was executed $20$ times, and the median result was reported for every triplet of sources, targets, and method.
 Both the Proteo version [^5] https://lorca.act.uji.es/gitlab/martini/malleability_benchmark/-/tree/Paper-ParallelSpawn and the experimental results [15] are publicly available.
-Proteo was used to execute the different approaches.
-Before reconfiguration, 5 iterations of Pi computation with MPI_Allgather were performed to ensure MPI initialization.
-MaM was configured with Baseline (B) and Merge (M), each with both strategies described in 4, or Merge without strategies.
-For shrinking, only Baseline plus the described techniques was compared against a stand-alone Merge.
-Each configuration was executed 20 times, and the median result was reported for every triplet of sources, targets, and method.
-5.2 Homogenous System Evaluation
-On the homogeneous cluster MNV, both strategies described in Section 4 were evaluated.
+5.2 Evaluation on Homogeneous Job Allocation
+Both spawning strategies described in Section 4 were evaluated on the homogeneous cluster MN5.
 In total, $5$ configurations were tested for expansion and $3$ for shrinking.
-Each configuration involved a single reconfiguration from $NS$ to $NT$ processes, using up to 42 combinations of nodes from the set 1,2,4,8,16,24,32{1,2,4,8,16,24,32}.
-Figure 4 reports the median reconfiguration times depending on the number of sources ($NS$) and targets ($NT$).
-The top Figure 4(a) shows expansion results, where the Merge method outperforms all others in 17 out of 21 cases ($80.9\%$).
-Parallel Merge follows closely, requiring up to $1.13\times$ more time due to additional synchronizations among groups.
+Each configuration involved a single reconfiguration from NS to NT processes, using up to $42$ nodes combinations from the set $\{1,2,4,8,16,24,32\}$.
+Figure 4 reports the median reconfiguration times depending on the number of initial nodes ($I$) and resulting nodes ($N$).
+Figure 4(a) (top) shows expansion results, where the Merge method outperforms all others in $17$ out of $21$ cases ($80.9\%$).
+Parallel Merge (in both versions) follows closely, requiring up to $1.13\times$ more time due to additional synchronizations among groups.
 This overhead grows when more than $8$ groups are created and the number of groups is not a power of two.
@@
 Parallel Baseline methods are consistently slower (up to $1.73\times$), except in one case ($1$ node to $16$ nodes).
-This overhead stems from spawning extra processes, which in addition leads to oversubscription.
-The bottom Figure 4(b) shows shrinking results.
-Shrinking without spawning processes is clearly advantageous, completing with a speedup of $1387\times$ over the other evaluated approaches.
+This overhead stems from spawning extra processes, which also leads to oversubscription.
+Figure 4(b) (bottom) shows shrinking results.
+In this case, avoiding process spawning is clearly advantageous, resulting in a speedup of $1387\times$ compared to the other evaluated approaches.
 These findings are consistent with previous work [17].
@@
 (b) Shrinkage.
-Figure 4: Resize times in homogenous system.
-Figure 5 summarizes the best method for each ($NS$, NTNT) pair based on statistical tests.
-Axes are defined with NS on the vertical and NT on the horizontal; thus, the upper triangle corresponds to expansion and the lower triangle to shrinking.
+Figure 4: Resizing times in an homogeneous allocation for MN5.
+The number of processes can be computed by multiplying the nodes by $112$ cores.
+Letter B refers to the Baseline method, M to Merge method and TS to the Merge shrink variant that terminates ranks.
+Figure 5 summarizes the best method for each ($I$, NN) pair based on statistical tests.
+Axes are defined with Initial Nodes on the vertical and Resulting Nodes on the horizontal; thus, the upper triangle corresponds to expansion and the lower triangle to shrinkage.
 When multiple methods appear in a cell, they are statistically equivalent, ordered by ascending time.
-The main difference arises in expansions with 8 or fewer groups, requiring at most 3 steps.
+The main difference arises in expansions with $8$ or fewer groups, requiring at most $3$ steps for the binary connection.
 In such cases, the parallel methods are preferred as they enable shrinking without spawning processes, while maintaining times comparable to Merge, though statistically higher.
-Figure 5: Median reconfiguration times in MNV.
-5.3 Improved Method Evaluation
-On the heterogeneous cluster NASP, only the improved strategy from Section 4.2 was evaluated.
-Three configurations were tested for expansion and two for shrinking.
-Each configuration involved a single reconfiguration from $NS$ to $NT$ processes, using up to 72 combinations of nodes from the set 1,2,4,6,8,10,12,14,16{1,2,4,6,8,10,12,14,16}.
+Figure 5: Preferred methods for process management phase depending on the number of nodes $I$ and NN, in MN5. The number of processes can be computed by multiplying the nodes by $112$ cores.
+5.3 Evaluation of Iterative Diffusive strategy
+The improved strategy described in Section 4.2 was evaluated only on the heterogeneous cluster NASP.
+Three configurations were tested for expansion and two for shrinkage.
+In this case, the Hypercube strategy is not included in the evaluation because it is unable to correctly spawn the processes.
+Each configuration involved a single reconfiguration from NS to NT processes, using up to $72$ combinations of nodes from the set $\{1,2,4,6,8,10,12,14,16\}$.
 In all cases, the number of nodes of each type was balanced (half of one type and half of the other).
 When only one node was used, the 20-core node was selected.
-Figure 6 shows the median reconfiguration times as a function of $NS$ and $NT$.
-The top Figure 6(a) reports expansion results, where the Merge method consistently outperforms the others.
-The iterative diffusive variant of Merge achieves comparable results, with at most a $1.25\times$ increase, but its cost grows with the number of generated groups.
-Baseline methods are always more expensive, ranging from $1.76\times$ to $4.92\times$ slower, due to oversubscription caused by spawning more processes.
-The bottom Figure 6(b) shows shrinking results.
-As in the previous evaluation, Merge is the preferred method with a minimum speedup of $20\times$.
-To ensure it can be applied, a parallel spawn is required.
-This approach also benefits future shrink operations, adding only a minor overhead during expansion.
-However, in difference to the previous evaluation, in 28 out of 32 cases ($87.5\%$), standalone Merge delivered statistically lower times than its iterative diffusive variant.
+Figure 6 shows the median reconfiguration times as a function of initial nodes ($I$) and resulting nodes ($N$).
+Figure 6(a) (top) reports expansion results, where the Merge method consistently outperforms the others.
+The iterative diffusive variant of Merge achieves comparable results, with at most a $1.25\times$ increase, but its cost increases slightly with the number of generated groups.
+Baseline methods are always more expensive, ranging from $1.76\times$ to $4.92\times$ slower, due to oversubscription caused by spawning additional processes.
+Figure 6(b) (bottom) presents the shrinking results.
+As observed in the expansion analysis, the B-based strategies are always the least efficient, as in this case M + TS method achieves speedups of at least $20\times$.
+To reach these benefits a parallel spawn must be performed beforehand, since TS requires that each MCW to be contained within a single node.
+Although this parallel spawning for expansions introduces an overhead compared to the previously optimal spawning method [17], the gains in subsequent shrinking operations outweigh that cost.
+Furthermore, in $28$ out of $32$ cases ($87.5\%$) for expanding and shrinking, Merge delivered statistically lower times than the iterative diffusive strategy.
 (a) Expansion.
 (b) Shrinkage.
-Figure 6: Resize times in heterogenous system.
+Figure 6: Resizing times in an heterogeneous allocation for NASP.
+The number of processes can be computed by multiplying the nodes by $52$ cores, with the exception for the node, where there are 2020 cores.
+Letter B refers to the Baseline method, M to Merge method and TS to the Merge shrink variant that terminates ranks.
 6 Conclusions and Future Work
-This work has introduced and evaluated two new strategies for spawning processes in the traditional MPI.
-Both approaches rely on parallel spawning with two main goals: (1) reducing the cost of process creation, and (2) enabling efficient shrink operations with the Merge method, the fastest available option.
-Achieving the second goal requires for each MCW to be fully contained within a node.
-Previously, this was not possible: when shrinking, processes that should be terminated remained as zombies because other ranks of the same MCW were still active on different nodes, preventing node reuse.
-The proposed strategies resolve this issue, enabling fast shrink operations.
-Between the two strategies, the Hypercube approach is limited to homogeneous, non-shared systems, while the Iterative Diffusive strategy also applies to heterogeneous or shared environments, at the cost of additional computation and memory overhead.
-Experimental results on both homogeneous and heterogeneous systems show that, in the worst case, the parallel approaches require up to $1.25\times$ the time of the previous best expansion technique, while enabling shrink operations that are at least $20\times$ faster.
+This work introduces and evaluates two new strategies for spawning processes in the traditional MPI.
+Both strategies rely on parallel spawning and have two main goals: (1) reducing the cost of process creation, and (2) enabling efficient shrink operations using the Merge method, the fastest available option.
+To achieve the second goal, each MCW must be fully contained within a single node.
+Previous Merge implementations could not achieve this because processes that should have been terminated during shrinkage remained as zombies preventing node reuse.
+Between the two strategies, the Hypercube approach is limited to homogeneous allocations, whereas the Iterative Diffusive strategy can also be used for resource allocations that use nodes with an heterogeneous number of cores.
+Experimental results on both homogeneous and heterogeneous allocations on different systems show that, in the worst case, the expansion with parallel approaches require up to $1.25\times$ times against the Merge method, but they enable shrink operations with a TS method that is at least $20\times$ faster than the Baseline method.
 In summary, this work provides a practical solution for enabling fast shrink operations without compromising expansion performance.
-Future work includes study how to further reduce the synchronization (Subsection 4.3) and connection (Subsection 4.4) overheads, with the aim of outperforming even the basic Merge in expansions.
-Another promising direction is the design of data redistribution strategies that minimize transfers, ensuring that processes after resizing retain as much of their data as possible.
+Future work includes studying how to further reduce the synchronization and connection (Subsections 4.3 and 4.4) overheads, with the aim of outperforming even the basic Merge expansions.
+Additionally, the load balancing challenges associated with heterogeneous hardware have not been considered in this work and must be addressed.
+Another promising direction is the design of data redistribution strategies that minimize transfers, ensuring that processes retain as much data as possible after resizing.
+Acknowledgements
+We would like thank Sergio Iserte for his valuable contributions to this work, his insightful discussions on possible improvements of the model, as well as his support in gaining access to the MN5 supercomputing facilities, which enabled the large-scale simulations presented in this article.
 Declaration of competing interest
 The authors have no conflicts of interest to declare that are relevant to the content of this article.
+Data availability
+The source codes for MaM are publicly available (https://lorca.act.uji.es/gitlab/martini/malleability_benchmark/-/tree/Paper-ParallelSpawn).
+All data used to create tables and figures in the Results section are available at [15].
 Funding sources
@@
 Martín-Álvarez was supported by the predoctoral fellowship ACIF/2021/260 from Valencian Region Government and European Social Funds.
+CRediT authorship contribution statement
+  • Iker Martín-Álvarez: Conceptualization, Methodology, Software, Investigation, Validation, Formal analysis, Data Curation, Visualization, Writing - Original Draft.
+  • José I. Aliaga: Methodology, Writing - Review & Editing, Supervision, Resources, Funding acquisition.
+  • Maribel Castillo: Methodology, Formal analysis, Writing - Review & Editing, Supervision.
+Declaration of generative AI use
+During the preparation of this work the authors used ChatGPT exclusively to refine the language and improve textual clarity.
+After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.